### v2

# The 2-systole on compact Kähler surfaces with positive scalar curvature

Abstract.
We study the 2-systole on compact Kähler surfaces of positive scalar curvature. For any such surface $(X,\omega)$, we prove the sharp estimate $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 12\pi$, with equality if and only if $X=\mathbb{P}^{2}$ and $\omega$ is the Fubini–Study metric. Using the classification of positive scalar curvature Kähler surfaces, we determine the optimal constant in each case and describe the corresponding rigid models. When $X$ is a non-rational ruled surface, we also give an independent analytic proof, adapting Stern’s level set method to the holomorphic fibration in Kähler setting.
1. Introduction
Systolic geometry studies quantitative lower bounds for nontrivial homology classes represented by cycles of minimal volume. Following Berger’s terminology [4, 3, 5] (see also [6]) and Gromov’s subsequent development of systolic geometry [12], if $(M,g)$ is a closed Riemannian manifold of dimension $n\geq k$, the kk–systole of $(M,g)$ is defined by
$\operatorname{sys}_{k}(g):=\inf\bigl\{\operatorname{Vol}_{g}(Z)\ \big|\ Z\text{ is a }k\text{-cycle with }[Z]\neq 0\in H_{k}(M;\mathbb{Z})\bigr\}.$
Here an integral $k$-cycle may be understood, for instance, as an integral current in the sense of geometric measure theory, and $\operatorname{Vol}_{g}(Z)$ denotes its $k$-dimensional mass with respect to $g$.
In the setting of positive scalar curvature (PSC), a particularly natural object is the 2-systole. Indeed, scalar curvature is obtained by averaging sectional curvatures over $2$-planes, and a basic mechanism for producing PSC metrics is to insert an $S^{2}$-component: for product metrics, one has $\text{scal}(g\oplus h)=\text{scal}(g)+\text{scal}(h)$, so $S^{2}(r)\times(N,h)$ carries PSC metrics once $r$ is sufficiently small, to a large extent, independently of the geometry of $N$. This suggests that, in many PSC constructions, the relevant geometric scale is encoded by a $2$-sphere direction, and hence by a $2$-systolic quantity.
In dimension three, the interplay between $2$-systole and PSC is by now well understood. If $(M^{3},g)$ is PSC, Schoen-Yau [21] proved that any area–minimizing surface in $M$ is homeomorphic to either $S^{2}$ or $\mathbb{RP}^{2}$, showing that area-minimizing surfaces in the PSC setting are forced to have spherical topology. Building on this, Bray-Brendle-Neves [7] established a sharp $\pi_{2}$-systolic inequality. Denote by $\operatorname{sys}^{\pi_{2}}_{2}(g)$ the infimum of the areas of homotopically nontrivial 22–spheres in $(M^{3},g)$. Then
$\min_{M}\text{scal}(g)\cdot\operatorname{sys}^{\pi_{2}}_{2}(g)\;\leq 8\pi,$ (1.1)
with equality if and only if the universal cover of $(M^{3},g)$ is isometric to the Riemannian product $S^{2}\times\mathbb{R}$ endowed with the round metric on $S^{2}$ and the flat metric on $\mathbb{R}$. In [22], Stern gave a new proof of (1.1) leveraging the level set method. The bound (1.1) was recently refined by a quantitative gap theorem of Xu [23], and there has been substantial further progress on systolic inequalities under scalar curvature assumptions; see for instance [26, 20, 18].
In real dimension four, by contrast, our current understanding of $2$-systoles under PSC assumptions is much more limited: even on $S^{2}\times S^{2}$ with positive scalar curvature, global upper bounds for $\operatorname{sys}_{2}(g)$ are only known under additional geometric hypotheses. In this paper we initiate a systematic study of $2$-systolic inequalities in dimension four under the extra assumption that the metric is Kähler. More precisely, given a compact Kähler surface $(X,\omega)$ we denote by $g_{\omega}$ the associated Riemannian metric and we write
$\operatorname{sys}_{2}(\omega):=\operatorname{sys}_{2}(g_{\omega}).$
We also write $S(\omega)$ for the Chern scalar curvature of $\omega$; for Kähler metrics, one has $2S(\omega)=\text{scal}(g_{\omega})$), so positivity of $S(\omega)$ is equivalent to positive Riemannian scalar curvature. With this convention, all our main inequalities are stated purely in terms of the Kähler form $\omega$.
Theorem 1.1.
Let $(X,\omega)$ be a compact Kähler surface with $S(\omega)>0$. Then
$\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\;\leq\;12\pi.$
Moreover, equality holds if and only if $X\cong\mathbb{P}^{2}$ and $\omega$ is the Fubini–Study metric.
The birational classification of PSC Kähler surfaces is by now complete. Yau proved that a compact Kähler surface with positive total scalar curvature has Kodaira dimension -$\infty$ [25]. By the Enriques–Kodaira classification, a minimal such surface is therefore either $\mathbb{P}^{2}$ or ruled [2, Chapter V]. In addition, LeBrun showed that for minimal complex surfaces of Kähler type, the existence of a PSC Riemannian metric is equivalent to the existence of a PSC Kähler metric, and this occurs precisely for $\mathbb{P}^{2}$ and ruled surfaces [17, Theorem 4]. The remaining birational step was completed by Brown, who proved that blow-ups preserve the sign of scalar curvature for sufficiently small data [8, Theorem B]. Thus a compact Kähler surface admits a PSC Kähler metric if and only if it is birational to $\mathbb{P}^{2}$ or to a ruled surface. It is therefore natural to restate Theorem 1.1 in a form that records how the optimal constant depends on the classification. This leads to the following three-way refinement.
Theorem 1.2.
Let $(X,\omega)$ be a compact Kähler surface with positive scalar curvature. Then:
  1. (1)
    (Theorem 3.4) If $X$ is birational to $\mathbb{P}^{2}$, then
    $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\;\leq\;12\pi,$
    with equality if and only if $(X,\omega)\cong(\mathbb{P}^{2},\omega_{\mathrm{FS}})$.
  2. (2)
    (Theorem 4.11) If $X$ is a rational ruled surface fibred over $\mathbb{P}^{1}$, then
    $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\;\leq\;8\pi,$
    with equality if and only if $X\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $\omega$ is the product Fubini–Study metric on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ (up to scaling).
  3. (3)
    (Theorems 4.5 and 5.4) If $X$ is a non-rational ruled surface fibred over a base curve $B$ of genus $g(B)\geq 1$, then
    $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\;\leq\;4\pi,$
    with equality if and only if $B$ is an elliptic curve, the universal cover of $X$ is biholomorphic to $\mathbb{P}^{1}\times\mathbb{C}$, and $\omega$ is induced by the product of the Fubini–Study metric on $\mathbb{P}^{1}$ and a flat metric on $\mathbb{C}$, with $\operatorname{sys}_{2}(\omega)$ is realized by a $\mathbb{P}^{1}$–fibre.
Idea of the proof
A key feature of the Kähler setting is that several quantities relevant to our problem admit a direct cohomological interpretation. In real dimension four, complex curves are calibrated by the Kähler form: if $(X,\omega)$ is a compact Kähler surface and $C\subset X$ is a smooth complex curve, then
$\operatorname{Area}_{\omega}(C)=\int_{C}\omega=[\omega]\cdot[C].$
More generally, if $C=\sum_{i}m_{i}C_{i}$ is an effective (possibly singular or reducible) curve, viewed as a positive integral real 22–cycle, then
$\omega]\cdot[C]=\int_{C}\omega=\sum_{i}m_{i}\int_{C_{i}^{\mathrm{reg}}}\omega.$
Thus the $\omega$–area of an effective curve is encoded entirely by the Kähler class $\omega]$ and the corresponding curve class in homology. This motivates the following algebraic analogue of the 22–systole.
Definition 1.3 (Holomorphic 22–systole).
Let $(X,\omega)$ be a compact Kähler manifold. The holomorphic 22–systole of a Kähler class $\omega]$ is defined by
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega]):=\inf\Bigl\{[\omega]\!\cdot\![C]\ \Big|\ C\subset X\text{ an effective curve},\ 0\neq[C]\in H_{2}(X;\mathbb{Z})\Bigr\}.$ (1.2)
In general, the infimum in (1.2) need not be taken over a nonempty set, since a compact Kähler manifold may contain no effective curves. In the PSC Kähler setting, however, this issue does not arise. By [24, Theorem 1.3], a compact complex manifold admits a Hermitian metric with positive Chern scalar curvature if and only if its canonical bundle is not pseudo-effective. If moreover the manifold is Kähler, then by [19, Theorem 1.1] this is equivalent to uniruledness. In particular, every compact PSC Kähler manifold contains plenty of effective curves, so $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$ is well-defined and finite in the setting of this paper.
We then introduce the scale-invariant functional
$\mathcal{J}_{X}([\omega]):=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\hat{S}([\omega]),$
where $\hat{S}([\omega])$ denotes the normalized average Chern scalar curvature of the Kähler class $\omega]$. Since this paper concerns Kähler surfaces, we may write
$\mathcal{J}_{X}([\omega])=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\frac{4\pi\,c_{1}(X)\cdot[\omega]}{[\omega]^{2}}.$
The importance of $\mathcal{J}_{X}$ comes from the elementary comparison
$\operatorname{sys}_{2}(\omega)\leq\operatorname{sys}^{\operatorname{hol}}_{2}([\omega]),$
because every effective curve gives a nontrivial real 22–cycle, together with the inequality
$\min_{X}S(\omega)\leq\hat{S}([\omega]).$
Hence
$\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq\mathcal{J}_{X}([\omega]).$
Therefore, the PSC $2$-systolic inequalities in this paper follow from sharp upper bounds for $\mathcal{J}_{X}$. Moreover, equality in the above comparison can occur only if $\omega$ has constant scalar curvature and $\operatorname{sys}_{2}(\omega)$ is realized by a holomorphic curve.
For PSC Kähler surfaces, this becomes a problem in intersection theory. Indeed, such surfaces are projective, and the Kähler cone admits a numerical description via the Nakai–Moishezon criterion. We may therefore regard $\mathcal{J}_{X}$ as a function on the cone of classes $\omega]$ satisfying
$\omega]^{2}>0,\qquad[\omega]\cdot C>0\ \text{for every irreducible curve }C\subset X,\qquad c_{1}(X)\cdot[\omega]>0.$
A first step, carried out in Section 2, is to show that $\mathcal{J}_{X}$ remains uniformly bounded on this region.
The main difficulty appears on blow-ups. If $\pi\colon X_{k}\to X$ is the blow-up of a minimal surface $X$ at $k$ points, then any Kähler class on $X_{k}$ can be written as
$\omega_{t}]=\pi^{*}[\omega]-\sum_{i=1}^{k}t_{i}E_{i},$
where $E_{i}$ are the exceptional divisors. In these coordinates, the averaged scalar curvature in $\mathcal{J}_{X_{k}}$ is determined cohomologically, whereas the holomorphic $2$-systole $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])$ depends a priori on the effective curve classes on $X_{k}$. Thus, for a fixed blow-up type, the configuration dependence of $\mathcal{J}_{X_{k}}$ is encoded in the Mori cone. In particular, special configurations may produce additional effective curves, hence additional inequalities cutting out the admissible region in the Kähler cone. From the point of view of upper bounds for $\mathcal{J}_{X_{k}}$, this can only make the problem more restrictive. Accordingly, the largest possible supremum should be expected in the most generic configuration, when the blow-up points are as general as possible.
The difficulty is that the quantity $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])$ depends a priori on all effective curve classes on $X_{k}$, and when the Mori cone is not rational polyhedral this involves infinitely many possibilities. Our key observation is that one can nevertheless control $\mathcal{J}_{X_{k}}([\omega_{t}])$ using only finitely many numerical parameters extracted from the coefficients $t_{i}$, together with a mass-shift argument that identifies the worst possible distribution of these coefficients under fixed coarse data. This leads to an explicit finite-dimensional optimization problem over a coarse admissible region determined by finitely many effective curves. In this way, the geometric problem is reduced to a manageable numerical one. We carry this out first for blow-ups of $\mathbb{P}^{2}$, and then for blow-ups of ruled surfaces.
For non-rational ruled surfaces, we also give an independent analytic proof, inspired by Stern’s level-set method [22]. In the Kähler setting, the holomorphic fibration provides a natural replacement for the harmonic map used in the Riemannian argument. This yields the same sharp constant and the same rigidity statement as the algebro-geometric approach.
Why the Kähler setting is special
It is useful to contrast the above discussion with the purely Riemannian PSC setting in higher dimensions. On a closed Riemannian manifold $(M^{n},g)$, the homological $k$-systole is only defined when $H_{k}(M;\mathbb{Z})\neq 0$, and even in that case a lower bound on scalar curvature does not in general force any upper bound for $\operatorname{sys}_{k}(g)$. A basic family of examples is
$M=S^{k}\times S^{n-k},\qquad g_{t}=t^{2}g_{S^{k}}\oplus g_{S^{n-k}},\qquad t\geq 1,$
where $g_{S^{m}}$ is the round metric of sectional curvature $1$. Then $\text{scal}(g_{t})$ stays uniformly positive, whereas the $k$-dimensional factor is stretched and $\operatorname{sys}_{k}(g_{t})\to\infty$ as $t\to\infty$. Thus positive scalar curvature alone does not control homological systoles in general.
By contrast, Kähler geometry provides additional structure. In basic Kähler examples, such as projective space, the relevant even-dimensional homology groups are already nontrivial, so the corresponding systolic questions are non-vacuous. More importantly for the present paper, on a compact Kähler surface the natural $2$-dimensional competitors are complex curves, and their areas are encoded cohomologically by the Kähler class.
In the PSC Kähler setting, the existence of such curves is moreover forced by the geometry itself: positive scalar curvature implies uniruledness, and hence the presence of many rational curves. This is the key difference from the general Riemannian situation. For PSC Kähler surfaces, one has canonical $2$-dimensional representatives and a rigid intersection-theoretic way to measure their areas. This is precisely what makes the holomorphic $2$-systole and the functional $\mathcal{J}_{X}$ the right objects for our problem.
Organization of the paper
In Section 2 we review the basic notions used throughout the paper. We also show that, for projective Kähler surfaces with $c_{1}(X)\cdot[\omega]>0$, the functional $\mathcal{J}_{X}$ is well-defined and satisfies $\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}<\infty$.
Section 3 is devoted to the case where the minimal model of $X$ is $\mathbb{P}^{2}$. We apply the mass-shift argument to obtain an upper bound for $\mathcal{J}_{\text{Bl}_{k}\mathbb{P}^{2}}$. This yields the sharp bound $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 12\pi$, together with the rigid model $(\mathbb{P}^{2},\omega_{\text{FS}})$.
In Section 4 we study ruled surfaces from an algebro-geometric perspective. We first treat the non-rational ruled case, which is technically simpler in our approach, and show that $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 4\pi$, together with the corresponding rigid model. We then recall the geometry of Hirzebruch surfaces and their blow-ups, and adapt the mass-shift argument to the rational ruled case. This leads to the sharp bound $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 8\pi$ for all PSC Kähler surfaces fibred over $\mathbb{P}^{1}$, with rigidity characterized by the product Fubini-Study metric on $\mathbb{P}^{1}\times\mathbb{P}^{1}$.
Section 5 provides an alternative proof in the non-rational ruled case. We introduce the level-set method for holomorphic fibrations $f\colon X\to B$ when the base curve $B$ admits a metric of non-positive Gaussian curvature.
Acknowledgment.
The author is deeply grateful to his advisor, Professor Gérard Besson, for introducing him to the study of positive scalar curvature and for his constant encouragement. The author also thanks Jian Wang for valuable comments on an earlier version of the manuscript.
2. Preliminaries
Let $(X^{n},\omega)$ be a compact Kähler manifold of complex dimension $n$. Locally we write
$\omega=\sqrt{-1}\,g_{i\bar{j}}\,dz_{i}\wedge d\bar{z}_{j}.$
We denote by
$H^{1,1}(X;\mathbb{R}):=H^{1,1}(X)\cap H^{2}(X;\mathbb{R})$
the space of real (1,1)(1,1)–classes, and by
$\mathcal{K}(X):=\{[\omega]\in H^{1,1}(X;\mathbb{R})\mid[\omega]>0\}$
the Kähler cone of $X$. It is an open convex cone in the finite-dimensional real vector space $H^{1,1}(X;\mathbb{R})$.
The (Chern–)Ricci form is defined by
$\operatorname{Ric}(\omega):=-\sqrt{-1}\,\partial\bar{\partial}\log\det(g_{i\bar{j}}).$
It is a real closed (1,1)(1,1)–form representing the first Chern class:
$c_{1}(X)=\frac{1}{2\pi}[\operatorname{Ric}(\omega)]\in H^{1,1}(X;\mathbb{R}).$
The (Chern–)scalar curvature $S(\omega)$ is then determined by
$S(\omega):=\frac{n\,\operatorname{Ric}(\omega)\wedge\omega^{n-1}}{\omega^{n}}.$
In particular, the normalized average scalar curvature
$\hat{S}([\omega]):=\frac{2n\pi\,c_{1}(X)\cup[\omega]^{n-1}}{[\omega]^{n}}$
depends only on the Kähler class $\omega]$. Clearly,
$\min_{X}S(\omega)\leq\hat{S}([\omega]),$
with equality if and only if $\omega$ is cscK.
Throughout the paper we write $g_{\omega}$ for the Riemannian metric associated to $\omega$, and denote by
$\operatorname{sys}_{2}(\omega):=\operatorname{sys}_{2}(g_{\omega})$
the $2$-systole. In contrast, the quantity that we will actually compute is the following class-dependent holomorphic $2$-systole:
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega]):=\inf\bigl\{[\omega]\cdot C\ \big|\ C\subset X\text{ an effective curve},\ [C]\neq 0\text{ in }H_{2}(X;\mathbb{Z})\bigr\}.$
Here an effective curve $C=\sum_{i}m_{i}C_{i}$ is viewed as an integral real 22–cycle, and we set
$\omega]\cdot C:=\sum_{i}m_{i}\int_{C_{i}^{\mathrm{reg}}}\omega.$
By abuse of notation, we do not distinguish a curve from its homology class when writing such pairings.
In particular, every effective curve determines a nontrivial integral 22–cycle, and therefore contributes a competitor in the definition of the $2$-systole. Hence we always have
$\operatorname{sys}_{2}(\omega)\leq\operatorname{sys}^{\operatorname{hol}}_{2}([\omega]).$ (2.1)
Note that $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$ depends only on the Kähler class $\omega]$, whereas $\operatorname{sys}_{2}(\omega)$ a priori depends on the specific metric in that class.
The basic scale-invariant functional that we will work with is the function
$\mathcal{J}_{X}:\mathcal{K}(X)\to\mathbb{R},\qquad\mathcal{J}_{X}([\omega]):=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\hat{S}([\omega]).$
Equivalently,
$\mathcal{J}_{X}([\omega])=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\frac{2n\pi\,c_{1}(X)\cup[\omega]^{n-1}}{[\omega]^{n}}.$
For Kähler surfaces ($n=2$) this becomes
$\mathcal{J}_{X}([\omega])=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\frac{4\pi\,c_{1}(X)\cdot[\omega]}{[\omega]^{2}}.$
Since $\operatorname{sys}^{\operatorname{hol}}_{2}(\lambda[\omega])=\lambda\,\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$ and $\hat{S}(\lambda[\omega])=\lambda^{-1}\hat{S}([\omega])$ for $\lambda>0$, the value of $\mathcal{J}_{X}([\omega])$ depends only on the ray $\mathbb{R}_{>0}\,[\omega]\subset\mathcal{K}(X)$.
In this paper we will only consider Kähler classes $\omega]$ with
$c_{1}(X)\cdot[\omega]>0,$
since these are the classes that may contain PSC Kähler metrics. We denote the corresponding subcone by
$\mathcal{K}^{+}(X):=\{[\omega]\in\mathcal{K}(X)\mid c_{1}(X)\cdot[\omega]>0\}.$
Lemma 2.1.
Let $V$ be a real vector space and let $K\subset V$ be an open cone. Let $f:K\to(0,\infty)$ and $J:K\to\mathbb{R}$ be functions satisfying
$f(\lambda x)=\lambda f(x),\qquad J(\lambda x)=J(x)$
for all $x\in K$ and all $\lambda>0$. Set
$K_{1}:=\{x\in K\mid f(x)=1\}.$
Assume that $K_{1}\neq\varnothing$. Then
$\sup_{x\in K}J(x)\;=\;\sup_{x\in K_{1}}J(x).$
Proof.
Define a map $\Phi:K\to K$ by
$\Phi(x):=\frac{x}{f(x)},\qquad x\in K.$
This is well-defined because $f(x)>0$ for every $x\in K$, and the conic property of $K$ implies $\Phi(x)\in K$. Moreover,
$f(\Phi(x))=f\!\left(\frac{x}{f(x)}\right)=\frac{1}{f(x)}\,f(x)=1,$
so that $\Phi(K)\subset K_{1}$.
Conversely, if $y\in K_{1}$, then $f(y)=1$ and $y=\Phi(y)$. Thus $\Phi(K)=K_{1}$. Finally, for every $x\in K$ we have
$J(\Phi(x))=J\!\left(\frac{x}{f(x)}\right)=J(x).$
Therefore
$\sup_{x\in K}J(x)=\sup_{x\in K}J(\Phi(x))=\sup_{y\in\Phi(K)}J(y)=\sup_{y\in K_{1}}J(y),$
which is the desired equality. ∎
Theorem 2.2.
Let $(X,\omega)$ be a projective Kähler surface with $\mathcal{K}^{+}(X)\neq\emptyset$. Then:
  1. (1)
    For every $\omega]\in\mathcal{K}^{+}(X)$ and $\omega\in[\omega]$ with positive scalar curvature, we have $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq\mathcal{J}_{X}([\omega])$, with equality if and only if $\omega$ is cscK and $\operatorname{sys}_{2}(\omega)$ is realized by a holomorphic $1$-cycle;
  2. (2)
    $\sup_{[\omega]\in\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])<+\infty$.
Proof.
Part (1) follows immediately from
$\min_{X}S(\omega)\leq\hat{S}([\omega])\quad\text{and}\quad\operatorname{sys}_{2}(\omega)\leq\operatorname{sys}^{\operatorname{hol}}_{2}([\omega]),$
with equality precisely when $\omega$ is cscK and $\operatorname{sys}_{2}(\omega)$ is realized by a holomorphic $1$-cycle. We therefore prove only Part (2).
Set $\mathcal{K}_{1}:=\{[\omega]\in\mathcal{K}^{+}(X)\mid\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=1\}$. By Lemma 2.1,
$\sup_{[\omega]\in\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])=\sup_{[\omega]\in\mathcal{K}_{1}}\mathcal{J}_{X}([\omega]).$
We first show that $\overline{\mathcal{K}_{1}}\subset\mathcal{K}(X)$. Assume for contradiction that there exists a sequence of Kähler classes $\omega_{\varepsilon}]\in\mathcal{K}_{1}$ such that $\omega_{\varepsilon}]\to[\omega_{0}]\in\partial\mathcal{K}(X)$.
Claim. There exists an effective curve $C\subset X$ such that
$\omega_{0}]\cdot C=0.$
Granting the claim, we obtain
$1=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{\varepsilon}])\leq[\omega_{\varepsilon}]\cdot C\to[\omega_{0}]\cdot C=0,$
a contradiction. Hence $\overline{\mathcal{K}_{1}}\subset\mathcal{K}(X)$.
Now fix a Euclidean norm $\|\cdot\|$ on the finite-dimensional real vector space $V:=H^{1,1}(X;\mathbb{R})$, and let $B:=\{\alpha\in V\mid\|\alpha\|=1\}$. Set $S_{1}:=\overline{\mathcal{K}_{1}}\cap B$. Then $S_{1}\subset\mathcal{K}(X)$ is compact. Consider the continuous function $F:\alpha\mapsto\alpha^{2}$ on $V$. Since $S_{1}\subset\mathcal{K}(X)$, we have $F>0$ on $S_{1}$, and therefore
$m:=\min_{S_{1}}F>0,\qquad M:=\max_{S_{1}}F<\infty.$
For any $\alpha\in\mathcal{K}_{1}$, define
$u:=\frac{\alpha}{\|\alpha\|}\in S_{1}.$
Then $\alpha^{2}=\|\alpha\|^{2}u^{2}$, so
$m\,\|\alpha\|^{2}\leq\alpha^{2}\leq M\,\|\alpha\|^{2}.$
Since $\alpha\mapsto c_{1}(X)\cdot\alpha$ is a continuous linear functional on $V$, there exists $C>0$ such that
$|c_{1}(X)\cdot\alpha|\leq C\|\alpha|\qquad\text{for all }\alpha\in V.$
Hence
$\left|\frac{c_{1}(X)\cdot\alpha}{\alpha^{2}}\right|\leq\frac{C\|\alpha\|}{m\|\alpha\|^{2}}=\frac{C^{\prime}}{\|\alpha\|}.$
It remains to show that $\|\alpha\|$ is bounded away from 0 on $\mathcal{K}_{1}$. Take any effective curve $F$. Since $\alpha\mapsto\alpha\cdot F$ is a continuous linear functional on $V$, there exists $\widetilde{C}>0$ such that
$|\alpha\cdot F|\leq\widetilde{C}\,\|\alpha\|\qquad\text{for all }\alpha\in V.$
On the other hand, for every $\alpha\in\mathcal{K}_{1}$, we have
$1=\operatorname{sys}^{\operatorname{hol}}_{2}(\alpha)\leq\alpha\cdot F.$
Therefore
$1\leq\alpha\cdot F\leq\widetilde{C}\,\|\alpha\|,$
so $\|\alpha\|\geq\widetilde{C}^{-1}$ on $\mathcal{K}_{1}$. This proves Part (2). ∎
Proof of the claim.
Set $\alpha_{0}:=[\omega_{0}]\in\partial\mathcal{K}(X)$. Since $\alpha_{0}$ is nef, we split into two cases.
Case 1: $\alpha_{0}^{2}>0$. If $\alpha_{0}\cdot C>0$ for every irreducible curve $C\subset X$, then $\alpha_{0}$ is ample by the Nakai–Moishezon criterion for surfaces, contradicting $\alpha_{0}\in\partial\mathcal{K}(X)$. Hence there exists an irreducible curve $C\subset X$ such that
$\alpha_{0}\cdot C=0.$
Case 2: $\alpha_{0}^{2}=0$. We first show that $c_{1}(X)\cdot\alpha_{0}=0$. Assume, for contradiction, that $c_{1}(X)\cdot\alpha_{0}>0$. Fix $t>3$ and set
$D_{t}:=K_{X}+t\alpha_{0}.$
We claim that $D_{t}$ is nef.
By the Mori’s Cone Theorem for smooth projective surfaces (see, for instance, [9, Theorem 7.49]), there exist countably many rational curves $C_{i}$ on $X$ such that
$0<-K_{X}\cdot C_{i}\leq 3,$
and
$\overline{\operatorname{NE}}(X)=\overline{\operatorname{NE}}(X)_{K_{X}\geq 0}+\sum_{i}\mathbb{R}_{\geq 0}[C_{i}],$
where
$\overline{\operatorname{NE}}(X)_{K_{X}\geq 0}:=\{\gamma\in\overline{\operatorname{NE}}(X)\mid K_{X}\cdot\gamma\geq 0\}.$
For $\gamma\in\overline{\operatorname{NE}}(X)_{K_{X}\geq 0}$ we have
$D_{t}\cdot\gamma=K_{X}\cdot\gamma+t\,\alpha_{0}\cdot\gamma\geq 0$
since $\alpha_{0}$ is nef. For a $K_{X}$-negative extremal curve $C_{i}$, we have
$D_{t}\cdot C_{i}=K_{X}\cdot C_{i}+t\,\alpha_{0}\cdot C_{i}\geq-3+t>0.$
Hence $D_{t}$ is nef.
Since both $D_{t}$ and $\alpha_{0}$ are nef, we have $D_{t}\cdot\alpha_{0}\geq 0$. On the other hand, $\alpha_{0}^{2}=0$ gives
$0\leq D_{t}\cdot\alpha_{0}=(K_{X}+t\alpha_{0})\cdot\alpha_{0}=K_{X}\cdot\alpha_{0}=-c_{1}(X)\cdot\alpha_{0}<0,$
which is a contradiction. Therefore, $c_{1}(X)\cdot\alpha_{0}=0$.
Since $\alpha_{0}\neq 0$ is nef with $\alpha_{0}^{2}=0$, the Hodge index theorem implies that the intersection form is non-positive on $\alpha_{0}^{\perp}$, and
$D_{t}\cdot\alpha_{0}=(K_{X}+t\alpha_{0})\cdot\alpha_{0}=K_{X}\cdot\alpha_{0}=-c_{1}(X)\cdot\alpha_{0}=0,$
we obtain $D_{t}^{2}\leq 0$. On the other hand, $D_{t}$ is nef, hence $D_{t}^{2}\geq 0$. Therefore $D_{t}^{2}=0$. The equality case in the Hodge index theorem now implies that $D_{t}$ is numerically proportional to $\alpha_{0}$, namely $D_{t}\equiv\lambda_{t}\,\alpha_{0}$ for some $\lambda_{t}\in\mathbb{R}$. Hence
$K_{X}\equiv a\,\alpha_{0}$
for some $a\in\mathbb{R}$. Since $\mathcal{K}^{+}(X)\neq\emptyset$, we may choose $\beta\in\mathcal{K}^{+}(X)$. Then
$(-K_{X})\cdot\beta=c_{1}(X)\cdot\beta>0,$
while $\alpha_{0}\cdot\beta>0$ because $\alpha_{0}$ is nef and nonzero. Therefore $a>0$. In particular, $X$ admits a Kähler class of positive total scalar curvature. By [15], $X$ is uniruled. Hence $\kappa(X)=$-$\infty$. By the Enriques–Kodaira classification of projective surfaces, the minimal model of $X$ is therefore either rational or ruled (see, for example, [2, Chapter V]).
If the minimal model of $X$ is ruled over a curve of genus $g$, then by the canonical divisor formula for ruled surfaces
$K_{X}^{2}=8(1-g)$
on the minimal model (see [14, Chapter V, §2]). Since each blow-up decreases $K^{2}$ by $1$, and since here
$K_{X}^{2}=(-K_{X})^{2}=a^{2}\alpha_{0}^{2}=0,$
it follows that necessarily $g=1$ and $X$ is already minimal. Thus the ruled case reduces to a minimal ruled surface over an elliptic curve. We now treat the two possibilities separately.
Subcase 2a: $X$ is rational. Since $X$ is rational, $\chi(\mathcal{O}_{X})=1$. Applying Riemann–Roch theorem to $D=-K_{X}$, we get
$\chi(\mathcal{O}_{X}(-K_{X}))=\chi(\mathcal{O}_{X})+\frac{(-K_{X})^{2}-K_{X}\cdot(-K_{X})}{2}=1+\frac{2K_{X}^{2}}{2}=1.$
By Serre duality,
$h^{2}(X,-K_{X})=h^{0}(X,2K_{X})=0,$
since $\kappa(X)=$-$\infty$. Hence
$h^{0}(X,-K_{X})-h^{1}(X,-K_{X})=1,$
and in particular
$h^{0}(X,-K_{X})\geq 1.$
Thus there exists a nonzero effective divisor $D\in|-K_{X}|$. Moreover,
$\alpha_{0}\cdot D=\frac{1}{a}(-K_{X})\cdot D=\frac{1}{a}(-K_{X})^{2}=0.$
Write $D=\sum_{j}m_{j}D_{j}$ as a sum of irreducible components. Since $\alpha_{0}$ is nef, each $\alpha_{0}\cdot D_{j}\geq 0$, and the equality $\alpha_{0}\cdot D=0$ forces $\alpha_{0}\cdot D_{j}=0$ for every component $D_{j}$. Thus any irreducible component of $D$ gives the desired curve.
Subcase 2b: $X$ is ruled over an elliptic curve. Since $X$ is a minimal ruled surface over an elliptic curve, let $C_{0}$ be a section of minimal self-intersection and $F$ be a fibre (as in Section 4). Then
$C_{0}^{2}=-e,\qquad e\geq-1,\qquad-K_{X}\equiv 2C_{0}+eF.$
Because $K_{X}$ is nef, we have
$0\leq(-K_{X})\cdot C_{0}=(2C_{0}+eF)\cdot C_{0}=2(-e)+e=-e.$
Hence $e=0$ or 1-1.
If $e=0$, then $K_{X}\equiv 2C_{0}\equiv a\,\alpha_{0}$, and therefore
$\alpha_{0}\cdot C_{0}=\frac{1}{a}(-K_{X})\cdot C_{0}=\frac{2}{a}\,C_{0}^{2}=0.$
Thus $C_{0}$ is the required irreducible effective curve.
If e=1e=-1, then $K_{X}\equiv 2C_{0}-F$. By [11, Proposition 3.2] (see also [16, Corollary 2.2]), there exists a smooth elliptic curve $E\subset X$ such that $E\equiv 2C_{0}-F$. Hence
$\alpha_{0}\cdot E=\frac{1}{a}(-K_{X})\cdot E=\frac{1}{a}(-K_{X})^{2}=0.$
Thus $E$ is the desired irreducible effective curve. This completes the proof in Case 2. ∎
Blow-ups will play a central role in what follows. Let $\pi\colon\tilde{X}\to X$ be the blow-up of $X$ at a point $p$, with exceptional divisor $E$. Given a Kähler class $\omega]$ on $X$, we consider the family of classes on $\tilde{X}$
$\omega_{t}]:=\pi^{*}[\omega]-t[E],\qquad t\geq 0.$
The Seshadri constant of $\omega]$ at $p$ is defined by
$\varepsilon([\omega];p):=\sup\{\,t\geq 0\mid\pi^{*}[\omega]-t[E]\ \text{is nef}\,\}=\inf_{C\ni p}\frac{[\omega]\cdot C}{\mathrm{mult}_{p}(C)},$
where the infimum is taken over irreducible curves $C\subset X$ passing through $p$, and $\mathrm{mult}_{p}(C)$ denotes the multiplicity of $C$ at $p$. In particular, for every $0<t<\varepsilon([\omega];p)$, the class $\omega_{t}]$ is Kähler on $\tilde{X}$.
Brown’s result [8, Theorem A] shows that if $\omega]$ contains a PSC Kähler metric, then for all sufficiently small $t>0$ the classes
$\omega_{t}]=\pi^{*}[\omega]-t[E]$
on the blow-up $\tilde{X}$ also contain PSC Kähler metrics. Therefore we may estimate $\mathcal{J}_{X}$ along such families and, ultimately, obtain uniform bounds for $\mathcal{J}_{X}$ on PSC Kähler classes arising from $\mathbb{P}^{2}$ and ruled surfaces under finitely many blow-ups.
3. The systolic inequality on $\mathbb{P}^{2}$ and its blow-up
In this section we study the $2$-systolic inequality on $\mathbb{P}^{2}$ and on its blow-up $\text{Bl}_{k}\mathbb{P}^{2}$ at $k$ points. On $\mathbb{P}^{2}$ we write $H$ for the hyperplane class, so that the Néron–Severi group is
$NS^{1}(\mathbb{P}^{2};\mathbb{R})=\langle H\rangle,\qquad H^{2}=1.$
Any Kähler class on $\mathbb{P}^{2}$ is of the form $\omega]=aH,\penalty 10000\ a>0$, and every effective curve class is a positive multiple of $H$. In particular, the holomorphic $2$-systole is
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=\inf_{C}[\omega]\cdot[C]=aH\cdot H=a,$
while
$c_{1}(\mathbb{P}^{2})=3H,\qquad c_{1}(\mathbb{P}^{2})\cdot[\omega]=3a,\qquad[\omega]^{2}=a^{2}.$
Hence
$\mathcal{J}_{\mathbb{P}^{2}}([\omega])=4\pi\,\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\,\frac{c_{1}(\mathbb{P}^{2})\cdot[\omega]}{[\omega]^{2}}=4\pi\cdot a\cdot\frac{3a}{a^{2}}=12\pi.$ (3.1)
We now pass to the blow-up $X_{k}:=\text{Bl}_{k}\mathbb{P}^{2}$ at $k$ points. The Néron–Severi group of $X_{k}$ is
$NS^{1}(X_{k};\mathbb{R})=\langle H,E_{1},\dots,E_{k}\rangle,$
where $H$ denotes the pullback of the hyperplane class and $E_{i}$ the exceptional divisors, with
$H^{2}=1,\qquad E_{i}^{2}=-1,\qquad H\cdot E_{i}=E_{i}\cdot E_{j}=0\ (i\neq j).$
In particular, any Kähler class can be written as
$\omega]=aH-\sum_{i=1}^{k}t_{i}E_{i},\qquad a,t_{i}\in\mathbb{R}_{>0},$
while a curve class in $H_{2}(X_{k};\mathbb{Z})$ is written
$C]=dH-\sum_{i=1}^{k}m_{i}E_{i},\qquad d,m_{i}\in\mathbb{Z},$
where we tacitly identify $H^{2}$ and $H_{2}$ via Poincaré duality.
In view of (3.1), our goal is to show that
$\mathcal{J}_{X_{k}}([\omega])<12\pi$
for every Kähler class $\omega]$ on $X_{k}$ and every $k\geq 1$. Since $\mathcal{J}_{X_{k}}$ is defined purely in terms of intersection numbers, the problem is intrinsically finite-dimensional. However, the complexity of the Mori cone $\overline{\mathrm{NE}}(X_{k})$ increases rapidly with $k$: for large $k$ there are many extremal rays, and for $k\geq 9$ (in the non–del Pezzo regime) the cone is not even finitely generated. A direct ray–by–ray analysis of all effective curve classes is therefore hopeless.
Instead of keeping track of each coefficient $t_{i}$ separately, it is convenient to work with the aggregate quantity
$S:=\sum_{i=1}^{k}t_{i}.$
Once a Kähler class $\omega]=aH-\sum t_{i}E_{i}$ is fixed, the numbers $t_{i}>0$ determine $S$, and the nef cone imposes a genuine geometric upper bound on $S$ which does not depend on the particular choice of coordinates. Moreover, on $X_{k}$ one has
$\omega]^{2}=a^{2}-\sum_{i=1}^{k}t_{i}^{2},\qquad c_{1}(X_{k})\cdot[\omega]=\left(3H-\sum_{i=1}^{k}E_{i}\right)\cdot\left(aH-\sum_{i=1}^{k}t_{i}E_{i}\right)=3a-\sum_{i=1}^{k}t_{i},$
so that, when $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$ is under control, $\mathcal{J}_{X_{k}}([\omega])$ depends on the tuple $(t_{1},\dots,t_{k})$ only through
$S=\sum_{i=1}^{k}t_{i},\qquad Q:=\sum_{i=1}^{k}t_{i}^{2}.$
For a given Kähler class we set $m:=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$. In particular, since each exceptional curve $E_{i}$ is effective and $\omega]\cdot E_{i}=t_{i}$, we always have $m\leq t_{i}$ for all $i$. Thus, if we fix $(a,m,S)$ and vary the individual $t_{i}$ subject to
$t_{i}\geq m,\qquad\sum_{i=1}^{k}t_{i}=S,$
then $a$ and $S$ are held fixed while $\mathcal{J}_{X_{k}}([\omega])$ changes only through the quantity $Q$ appearing in $\omega]^{2}$. In particular, for fixed $(a,m,S)$ the maximal possible value of $\mathcal{J}_{X_{k}}([\omega])$ is attained when $Q$ is as large as allowed by these constraints. This leads to the following elementary “mass–shift” property, which will play a crucial role not only in the present section but also in our later analysis of Hirzebruch surfaces, i.e. geometrically ruled surfaces over $\mathbb{P}^{1}$.
Proposition 3.1.
Let $k\geq 2$, $m>0$, and $S\geq km$. Consider
$\Omega=\Bigl\{t=(t_{1},\dots,t_{k})\in\mathbb{R}^{k}:\ t_{i}\geq m\ \text{for all }i,\ \sum_{i=1}^{k}t_{i}=S\Bigr\}.$
For $Q(t):=\sum_{i=1}^{k}t_{i}^{2}$ one has
$\sup_{t\in\Omega}Q(t)=(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2},$
and the supremum is attained exactly (up to permutation of coordinates) at
$t=(\underbrace{m,\dots,m}_{k-1},\,S-(k-1)m).$
Proof.
Since $S\geq km$, the point $(S/k,\dots,S/k)$ belongs to $\Omega$, so $\Omega\neq\emptyset$. In particular, $\Omega$ is compact in $\mathbb{R}^{k}$, and since $Q$ is continuous, $Q$ attains its maximum on $\Omega$. Let $t=(t_{1},\dots,t_{k})\in\Omega$ be a maximizer.
We first show that at most one coordinate of $t$ is strictly larger than $m$. Suppose by contradiction that there exist $i\neq j$ with
$t_{i}\geq t_{j}>m.$
Choose $\delta\in(0,\,t_{j}-m]$ and define
$t_{i}^{\prime}=t_{i}+\delta,\qquad t_{j}^{\prime}=t_{j}-\delta,\qquad t_{\ell}^{\prime}=t_{\ell}\quad(\ell\notin\{i,j\}).$
Then $t^{\prime}\in\Omega$, and
$Q(t^{\prime})-Q(t)$ $=(t_{i}+\delta)^{2}+(t_{j}-\delta)^{2}-(t_{i}^{2}+t_{j}^{2})$
$=2\delta(t_{i}-t_{j})+2\delta^{2}\;\geq 2\delta^{2}>0,$
which contradicts the maximality of $t$. Hence at most one coordinate of $t$ exceeds $m$.
Since all coordinates satisfy $t_{i}\geq m$, it follows that exactly k1k-1 coordinates are equal to $m$, and the remaining one equals S(k1)mS-(k-1)m. The condition $S\geq km$ ensures $S-(k-1)m\geq m$, so such a point lies in $\Omega$. Evaluating $Q$ there gives
$Q(t)=(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2}.$
Thus
$\sup_{s\in\Omega}Q(s)=Q(t)=(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2}.$
Finally, if $t\in\Omega$ satisfies
$Q(t)=(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2},$
then $t$ is a maximizer, and the above argument shows that (up to permutation) $t$ must have the stated form. The case $S=km$ corresponds to S(k1)m=mS-(k-1)m=m, i.e. $t_{i}=m$ for all $i$. ∎
In the numerical optimization below, we enlarge the geometric problem to a coarser numerical one. More precisely, we deliberately forget the full dependence of $m=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])$ on the vector $t=(t_{1},\dots,t_{k})$, and retain only the necessary inequalities forced by the effective curves $E_{i}$ and $H-E_{i}-E_{j}$. This produces a larger numerical feasible region, so any upper bound obtained there automatically yields an upper bound for the geometric supremum of $\mathcal{J}_{X_{k}}$. Namely, for every Kähler class $\omega_{t}$ we have
$m\leq\omega_{t}\!\cdot E_{i}=t_{i},\qquad m\leq\omega_{t}\!\cdot(H-E_{i}-E_{j})=a-t_{i}-t_{j},$
and in particular $0<m\leq a/3$ and $S=\sum_{i}t_{i}\geq km$. We then rewrite
$\mathcal{J}_{X_{k}}([\omega_{t}])=4\pi\,m\,\frac{3a-S}{a^{2}-Q}=:4\pi\,\phi_{k}(m,S,Q),\qquad Q=\sum_{i=1}^{k}t_{i}^{2}.$
Fix $(a,m,S)$ in the above coarse admissible region and consider all vectors $t=(t_{1},\dots,t_{k})$ satisfying only the numerical conditions
$t_{i}\geq m,\qquad\sum_{i=1}^{k}t_{i}=S,$
which are necessary for any geometric class with systole $m$ and sum $S$. Since the factor 3aS3a-S is fixed and $a^{2}-Q>0$ on the Kähler cone, the map $Q\mapsto m(3a-S)/(a^{2}-Q)$ is increasing in $Q$. Therefore, for any geometric vector $t$ with given $(m,S)$ we obtain the upper bound
$\phi_{k}(m,S,Q(t))\leq\phi_{k}\bigl(m,S,Q_{\max}(m,S)\bigr),$
where $Q_{\max}(m,S)$ denotes the maximal possible value of $\sum t_{i}^{2}$ under the constraints $t_{i}\geq m$ and $\sum t_{i}=S$.
By Proposition 3.1, this maximum is achieved at the extremal vector
$t^{*}=(m,\dots,m,\,S-(k-1)m),$
and hence
$Q_{\max}(m,S)=(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2}.$
Consequently, for every Kähler class $\omega_{t}$ we have the numerical estimate
$\mathcal{J}_{X_{k}}([\omega_{t}])\leq 4\pi\,m\,\frac{3a-S}{a^{2}-Q_{\max}(m,S)}.$
In particular, to bound $\sup\mathcal{J}_{X_{k}}$ from above, it suffices to maximize the right-hand side over the coarse admissible region in $(m,S)$ determined above, reducing the problem to a purely numerical optimization.
Lemma 3.2.
Let $X_{k}:=\operatorname{Bl}_{k}\mathbb{P}^{2}$, and let $\omega_{t}]=aH-\sum_{i=1}^{k}t_{i}E_{i}\in\mathcal{K}^{+}(X_{k})$, where $a>0$ and $k\geq 1$. Then
$\mathcal{J}_{X_{k}}([\omega_{t}])<\mathcal{J}_{\mathbb{P}^{2}}(aH)=12\pi.$
Proof.
We first treat the case $k=1$. In this situation the Kähler class is $\omega_{t}]=aH-tE$, and a direct computation (see Example 3.5) shows that
$\mathcal{J}_{\operatorname{Bl}_{1}\mathbb{P}^{2}}([\omega_{t}])=4\pi\,\min\{a-t,t\}\,\frac{3a-t}{a^{2}-t^{2}}\;\leq\;\frac{20\pi}{3}<12\pi.$
Thus, it remains to consider $k\geq 2$.
For $k\geq 2$, the exceptional curves $E_{i}$ and the strict transforms of lines through two points, with classes $H-E_{i}-E_{j}$ ($1\leq i\neq j\leq k$), are effective. Hence
$\omega_{t}\cdot E_{i}=t_{i}>0,\qquad\omega_{t}\cdot(H-E_{i}-E_{j})=a-t_{i}-t_{j}>0,$
are candidates for the holomorphic $2$-systole. Set $m:=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])\leq\min\{t_{i},\ a-t_{i}-t_{j}:\ 1\leq i\neq j\leq k\}$. Then
$\mathcal{J}_{X_{k}}([\omega_{t}])=4\pi\,m\,\frac{3a-S}{a^{2}-Q}.$
For convenience, we denote the corresponding numerical factor by
$\phi_{k}(t):=m\,\frac{3a-S}{a^{2}-Q},$
where $m=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])$, $S=\sum_{i}t_{i}$, and $Q=\sum_{i}t_{i}^{2}$. We now estimate $\phi_{k}(t)$ from above by optimizing over the enlarged numerical feasible region described above. Observe that $\mathcal{J}$ (and hence $\phi_{k}$) is invariant under rescaling of the Kähler class. Thus, without loss of generality, we may assume $a=1$ in what follows. With this normalization,
$\phi_{k}(t)=m\cdot\frac{3-S}{1-Q}.$
By definition of $m$ we have
$t_{i}\geq m,\qquad 1-t_{i}-t_{j}\geq m\quad(1\leq i\neq j\leq k),$
so in particular $0<m\leq 1/3$. Summing $t_{i}\geq m$ also gives $S\geq km$.
Fix $m>0$ and $S\geq km$, and consider all $t=(t_{1},\dots,t_{k})$ with
$t_{i}\geq m,\qquad\sum_{i=1}^{k}t_{i}=S.$
By Proposition 3.1, we have $Q\leq(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2}$, with equality at $t^{*}=(m,\dots,m,\,S-(k-1)m)$ up to permutation of coordinates. Since the denominator 1Q1-Q is decreasing in $Q$, we obtain
$\phi_{k}(t)=m\,\frac{3-S}{1-Q}\;\leq\;m\,\frac{3-S}{1-(k-1)m^{2}-\bigl(S-(k-1)m\bigr)^{2}}=:F(m,S),$ (3.2)
for every geometric vector $t$ arising from a class in $\mathcal{K}^{+}(X_{k})$. We now work on the admissible domain
$km\leq S\leq 1+(k-3)m,\quad 0<m\leq\tfrac{1}{3},\quad 1-(k-1)m^{2}-\bigl(S-(k-1)m\bigr)^{2}>0.$ (3.3)
We distinguish two parts.
Case 1: $2\leq k\leq 8$. Set B:=1+(k3)mB:=1+(k-3)m. We claim that for every $(m,S)$ in the admissible domain, $F(m,S)\leq F(m,B)$. Indeed, a direct computation gives
$F(m,B)-F(m,S)=\frac{(B-S)\,L_{k}(m,S)}{\bigl(4-(k+3)m\bigr)\bigl(1-(k-1)m^{2}-(S-(k-1)m)^{2}\bigr)},$
where $L_{k}(m,S)=S(2-(k-3)m)+k(k-1)m^{2}-3(k+1)m+2$.
Since $B-S\geq 0$, 4(k+3)m>04-(k+3)m>0, and $1-(k-1)m^{2}-(S-(k-1)m)^{2}>0$, it remains to show that $L_{k}(m,S)\geq 0$. For $2\leq k\leq 8$ and $0<m\leq 1/3$, we have
$2-(k-3)m\geq 2-\frac{k-3}{3}=\frac{9-k}{3}>0,$
so $L_{k}(m,S)$ is increasing in $S$. Hence
$L_{k}(m,S)\geq L_{k}(m,km)=2k\,m^{2}-(k+3)m+2.$
The discriminant of the quadratic polynomial $2k\,m^{2}-(k+3)m+2$ is
$(k+3)^{2}-16k=(k-1)(k-9)<0\qquad(2\leq k\leq 8),$
and its leading coefficient is positive. Therefore $2k\,m^{2}-(k+3)m+2>0$ for all $m$, and hence $L_{k}(m,S)>0$. This proves
$F(m,S)\leq F\bigl(m,1+(k-3)m\bigr).$
A direct computation yields
$F\bigl(m,1+(k-3)m\bigr)=\frac{2-(k-3)m}{\,4-(k+3)m\,}=:G_{k}(m).$ (3.4)
The feasibility condition $1-(k-1)m^{2}-(1-2m)^{2}>0$ is equivalent to
$0<m<\frac{4}{k+3}.$
Differentiating (3.4) gives
$G_{k}^{\prime}(m)=\frac{18-2k}{(4-(k+3)m)^{2}}>0\qquad(2\leq k\leq 8).$
Since $4/(k+3)\geq 1/3$ for $k\leq 8$, the constraint $m\leq 1/3$ is active, and thus
$\sup_{m}G_{k}(m)=G_{k}\Bigl(\frac{1}{3}\Bigr)=1.$
This value is attained at
$m=\frac{1}{3},\qquad S=1+(k-3)\frac{1}{3}=\frac{k}{3}.$
Hence
$\sup_{t}\phi_{k}(t)\leq 1\qquad\text{for }2\leq k\leq 8.$
Case 2: $k\geq 9$. We claim that
$F(m,S)\leq\frac{1}{2}$
throughout the admissible domain. Indeed, this inequality is equivalent to
$2m(3-S)\leq 1-(k-1)m^{2}-(S-(k-1)m)^{2}.$
Set y:=S(k1)my:=S-(k-1)m. Then the difference between the right-hand side and the left-hand side is
$1-(k-1)m^{2}-y^{2}-2m(3-(k-1)m-y)=1-(y-m)^{2}+km^{2}-6m.$
Since $km\leq S\leq 1+(k-3)m$, we have $m\leq y\leq 1-2m$, and therefore $0\leq y-m\leq 1-3m$. It follows that
$1-(y-m)^{2}+km^{2}-6m\geq 1-(1-3m)^{2}+km^{2}-6m=(k-9)m^{2}\geq 0.$
Thus
$F(m,S)\leq\frac{1}{2}\qquad\text{for all }k\geq 9.$
For $k=9$, equality is achieved whenever $S=1+6m$, equivalently y=12my=1-2m, so
$\sup_{t}\phi_{9}(t)\leq\frac{1}{2}.$
[…]
$\sup_{t}\phi_{k}(t)\leq\frac{1}{2}\qquad\text{for }k\geq 10.$
Collecting the two cases, we conclude that
$\sup_{t}\phi_{k}(t)\leq\begin{cases}1,&2\leq k\leq 8,\\[2.0pt] \frac{1}{2},&k\geq 9.\end{cases}$
In particular,
$\sup_{t}\phi_{k}(t)\leq 1\qquad\text{for all }k\geq 2.$
Combining this with the case $k=1$ treated at the beginning, we obtain
$\mathcal{J}_{\text{Bl}_{k}\mathbb{P}^{2}}([\omega_{t}])<\mathcal{J}_{\mathbb{P}^{2}}([\omega])=12\pi$
for all $k\geq 1$, as claimed. ∎
Remark 3.3.
The dependence of the upper bounds for $\mathcal{J}_{\operatorname{Bl}_{k}\mathbb{P}^{2}}$ on $k$ is in perfect agreement with the standard classification of the blow-ups $\operatorname{Bl}_{k}\mathbb{P}^{2}$ according to the positivity of the canonical class. Indeed,
$c_{1}^{2}(\operatorname{Bl}_{k}\mathbb{P}^{2})=\left(3H-\sum_{i=1}^{k}E_{i}\right)^{2}=9-k.$
Thus $k=9$ is exactly the borderline case where $c_{1}^{2}=0$, while $c_{1}^{2}>0$ for $k\leq 8$ and $c_{1}^{2}<0$ for $k\geq 10$. In the del Pezzo range $k\leq 8$ (for points in general position), this agrees with the usual positivity picture. We also note that for $k\geq 5$ additional (1)(-1)-curves appear, for instance conics of class $2H-\sum_{i\in I}E_{i}$ through five points. Their intersection with $\omega_{t}]$ gives further inequalities (such as $2a-\sum_{i\in I}t_{i}>0$) which can only shrink the feasible region for $(m,S)$ and hence potentially lower the true supremum of $\mathcal{J}_{X_{k}}$. Thus the upper bounds of $\mathcal{J}_{\text{Bl}_{k}\mathbb{P}^{2}}$ above should not be expected to be sharp in general.
As a consequence of the previous estimates, we obtain the desired global systolic inequality on $\mathbb{P}^{2}$ and its blow-ups.
Theorem 3.4.
Let $X_{0}=\mathbb{P}^{2}$ and let $X_{k}=\text{Bl}_{k}\mathbb{P}^{2}$ be the blow-up at $k$ points for $k\geq 1$. Suppose $\omega_{k}$ is a PSC Kähler metric on $X_{k}$. Then
$\min_{X_{k}}S(\omega_{k})\cdot\operatorname{sys}_{2}(\omega_{k})\;\leq\;12\pi,$ (3.5)
with equality if and only if $k=0$, in which case $\omega_{0}$ is the Fubini-Study metric (after normalization) and $\operatorname{sys}_{2}(\omega)$ is achieved by $\mathbb{P}^{1}$. In particular, (3.5) holds strictly for all blow-ups $X_{k}$ with $k\geq 1$.
Proof.
For $k=0$, equation (3.1) shows that $\mathcal{J}_{\mathbb{P}^{2}}([\omega])=12\pi$ for every Kähler class $\omega]$ on $\mathbb{P}^{2}$. Hence Theorem 2.2 (1) gives
$\min_{\mathbb{P}^{2}}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 12\pi.$
For $k\geq 1$, Lemma 3.2 shows that every class $\omega_{k}]\in\mathcal{K}^{+}(X_{k})$ satisfies
$\mathcal{J}_{X_{k}}([\omega_{k}])<12\pi.$
If $\omega_{k}$ is a PSC Kähler metric on $X_{k}$, then $\omega_{k}]\in\mathcal{K}^{+}(X_{k})$, so Theorem 2.2 (1) again yields
$\min_{X_{k}}S(\omega_{k})\cdot\operatorname{sys}_{2}(\omega_{k})\leq\mathcal{J}_{X_{k}}([\omega_{k}])<12\pi.$
This proves the strict inequality for all $k\geq 1$.
It remains to characterize the equality case for $k=0$. If
$\min_{\mathbb{P}^{2}}S(\omega_{0})\cdot\operatorname{sys}_{2}(\omega_{0})=12\pi,$
then equality must hold in Theorem 2.2 (1). Hence $\omega_{0}$ is cscK and $\operatorname{sys}_{2}(\omega_{0})$ is realized by a holomorphic $1$-cycle. Since every Kähler class on $\mathbb{P}^{2}$ is proportional to $H$, such a cscK metric is homothetic to the Fubini–Study metric, and the holomorphic $2$-systole is realized by a projective line $\mathbb{P}^{1}\subset\mathbb{P}^{2}$. Conversely, every metric homothetic to the Fubini–Study metric realizes equality. ∎
Example 3.5.
We can compute the precise supremum of $\mathcal{J}_{X_{k}}([\omega])$ when $k=1,2$. For $k=1$, any Kähler class can be written as $\omega]=aH-tE,$ for a>0,t>0,a>ta>0,\ t>0,\ a>t. Also, we have
$c_{1}(X_{1})=3H-E,\qquad c_{1}(X_{1})\cdot[\omega]=3a-t,\qquad[\omega]^{2}=a^{2}-t^{2}.$
The Mori cone of $X_{1}$ is generated by $E$ and HEH-E, so
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=\min\{[\omega]\cdot E,\,[\omega]\cdot(H-E)\}=\min\{t,\ a-t\}.$
Since $\mathcal{J}_{X_{1}}([\omega])$ is invariant under overall scaling of $\omega]$, it only depends on the ratio $x:=t/a\in(0,1)$. We may therefore assume $a=1$, and obtain
$\mathcal{J}_{X_{1}}([\omega])=4\pi\,\min\{t,1-t\}\,\frac{3-t}{1-t^{2}}=:h_{1}(t),\qquad 0<t<1.$
A direct computation shows that $h_{1}(t)$ is strictly increasing on $(0,\tfrac{1}{2}]$, so its maximum is attained at $t=\tfrac{1}{2}$, i.e. in the class proportional to $H-\tfrac{1}{2}E$. Hence
$\sup_{\mathcal{K}(X_{1})}\mathcal{J}_{X_{1}}([\omega])=\mathcal{J}_{X_{1}}\left([H-\tfrac{1}{2}E]\right)=\frac{20\pi}{3}.$
For $k=2$, any Kähler class can be written as $\omega]=aH-t_{1}E_{1}-t_{2}E_{2},\penalty 10000\ a>0,\ t_{i}>0$. And,
$c_{1}(X_{2})=3H-E_{1}-E_{2},\qquad c_{1}(X_{2})\cdot[\omega]=3a-t_{1}-t_{2},\qquad[\omega]^{2}=a^{2}-t_{1}^{2}-t_{2}^{2}.$
The Mori cone of $X_{2}$ is generated by the (1)(-1)-curves $E_{1},E_{2},H-E_{1}-E_{2}$, so
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=\min\bigl\{t_{1},\ t_{2},\ a-t_{1}-t_{2}\bigr\}.$
Again $\mathcal{J}_{X_{2}}([\omega])$ is invariant under overall scaling of $\omega]$, so we may set $a=1$ and obtain
$\mathcal{J}_{X_{2}}([\omega])=4\pi\,\min\bigl\{t_{1},\ t_{2},\ 1-t_{1}-t_{2}\bigr\}\frac{3-t_{1}-t_{2}}{1-t_{1}^{2}-t_{2}^{2}}=:h_{2}(t_{1},t_{2}),\qquad 0<t_{1}+t_{2}<1.$
A straightforward analysis of the resulting two-variable function shows that the maximum of $h_{2}$ occurs at $t_{1}=t_{2}=\tfrac{1}{3}$, so that
$\sup_{\mathcal{K}(X_{2})}\mathcal{J}_{X_{2}}([\omega])=\mathcal{J}_{X_{2}}\Bigl(\bigl[H-\tfrac{1}{3}E_{1}-\tfrac{1}{3}E_{2}\bigr]\Bigr)=4\pi.$
In both cases $k=1,2$, for any PSC Kähler metric $\omega$ on $X_{k}$ we have
$\min_{X_{k}}S(\omega)\cdot\operatorname{sys}_{2}(\omega)<\mathcal{J}_{X_{k}}([\omega])\leq\sup_{\mathcal{K}(X_{k})}\mathcal{J}_{X_{k}},$
and the inequality is strict since there is no cscK metric on $\operatorname{Bl}_{k}\mathbb{P}^{2}$ for $k=1,2$.
4. Systolic inequalities on ruled surfaces
In this section we study PSC Kähler metrics on ruled surfaces and establish a uniform upper bound for the $2$-systole. The following theorem summarizes the global picture.
Theorem 4.1.
Let $X\to B$ be a ruled surface (not necessarily minimal) fibred over a smooth complex curve $B$, and let $\omega$ be a PSC Kähler metric on $X$. Then
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)\;\leq\;8\pi.$
Moreover, equality holds if and only if $B\cong\mathbb{P}^{1}$ and $X\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$, and, up to overall scaling, $\omega$ is the product Fubini–Study metric.
The proof of Theorem 4.1 will be divided according to the genus of the base curve. In the non-rational case $g(B)\geq 1$ we actually obtain the sharper bound $4\pi$ (Theorem 4.5), whereas in the rational case $B\cong\mathbb{P}^{1}$ we show that the optimal constant is $8\pi$, with rigidity only on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ (Theorem 4.11).
4.1. Geometry of ruled surfaces and notation
Let $X_{0}\to B$ be a minimal ruled surface over a smooth complex curve $B$ of genus $g$. Let $F$ denote the fibre class, and let $C_{0}$ be a section of minimal self-intersection. We write
$F^{2}=0,\qquad C_{0}\cdot F=1,\qquad C_{0}^{2}=-e,$
where $e\in\mathbb{Z}$ is the invariant of the ruled surface. For $B\cong\mathbb{P}^{1}$, one always has $e\geq 0$, and $X_{0}\cong\mathbb{F}_{e}$. For $g\geq 2$, one also has $e\geq 0$. When $g=1$, besides the cases $e\geq 0$, there is one exceptional indecomposable minimal ruled surface with e=1e=-1.
In all cases, the first Chern class is $c_{1}(X_{0})=2C_{0}+(2-2g+e)F$. A numerical class on $X_{0}$ can be written as $\omega]=aC_{0}+bF$, and one computes
$\omega]^{2}=(aC_{0}+bF)^{2}=2ab-ea^{2},$
and
$c_{1}(X_{0})\cdot[\omega]=\bigl(2C_{0}+(2-2g+e)F\bigr)\cdot(aC_{0}+bF)=2b+(2-2g-e)a.$
When $e\geq 0$, the Kähler cone is given by $a>0,\penalty 10000\ b>ae$. In the exceptional elliptic case $g=1$, e=1e=-1, the Kähler cone is given by $a>0,\penalty 10000\ b>-\frac{a}{2}$. Indeed, in this case the extra effective class $2C_{0}-F=-K_{X_{0}}$ appears, and positivity on $F$ and $2C_{0}-F$ is equivalent to the above condition.
Proposition 4.2.
Let $X_{0}$ be a minimal ruled surface over a smooth complex curve $B$ of genus $g$, with invariant $e$. Then
$\sup_{\mathcal{K}(X_{0})}\mathcal{J}_{X_{0}}([\omega])=\begin{cases}\displaystyle 4\pi\,\dfrac{e+4}{e+2},&g=0,\\[8.0pt] \displaystyle 4\pi,&g\geq 1.\end{cases}$ (4.1)
Moreover:
  1. (1)
    when $g=0$, the supremum is attained precisely on classes proportional to $C_{0}+(e+1)F$;
  2. (2)
    when $g=1$ and $e\geq 0$, the supremum is attained precisely when $b/a\geq e+1$;
  3. (3)
    when $g=1$ and e=1e=-1, the supremum is attained precisely when $b/a\geq 0$;
  4. (4)
    when $g>1$, the supremum is not attained.
Proof.
We split the proof according to the value of $e$.
Case 1: $e\geq 0$. In this case it is well known that the Mori cone $\overline{\operatorname{NE}}(X_{0})$ is generated by the numerical classes of $F$ and $C_{0}$. Hence
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=\min\{[\omega]\cdot F,\,[\omega]\cdot C_{0}\}=\min\{a,\ b-ae\}.$
Therefore
$\mathcal{J}_{X_{0}}([\omega])=\min\{a,\ b-ae\}\cdot 4\pi\,\frac{2b+(2-2g-e)a}{2ab-ea^{2}}.$
Since $\mathcal{J}_{X_{0}}$ is homogeneous of degree 0 in $\omega]$, it depends only on the ratio $t:=b/a$. The Kähler condition becomes $t>e$, and
$\mathcal{J}_{X_{0}}([\omega])=4\pi\,\phi_{e,g}(t),\qquad\phi_{e,g}(t):=\min\{1,\ t-e\}\cdot\frac{2t+2-2g-e}{2t-e},\qquad t>e.$
We now distinguish according to $g$.
Subcase 1a: $g\geq 1$. Since $2-2g\leq 0$, we have $2t+2-2g-e\leq 2t-e$, and hence
$\frac{2t+2-2g-e}{2t-e}\leq 1.$
Therefore
$\phi_{e,g}(t)\leq\min\{1,t-e\}\leq 1,$
so
$\sup_{t>e}\phi_{e,g}(t)\leq 1.$
If $g=1$, then $\phi_{e,1}(t)=\min\{1,t-e\}$, so $\sup_{t>e}\phi_{e,1}(t)=1$, and the supremum is attained precisely when $t-e\geq 1$, i.e. $b/a\geq e+1$. If $g>1$, then 22g<02-2g<0, so $\phi_{e,g}(t)<\min\{1,t-e\}\leq 1$ for all $t>e$. On the other hand,
$\lim_{t\to+\infty}\phi_{e,g}(t)=\lim_{t\to+\infty}\frac{2t+2-2g-e}{2t-e}=1.$
Hence $\sup_{t>e}\phi_{e,g}(t)=1$, but it is not attained. Thus, for all $g\geq 1$ and $e\geq 0$,
$\sup_{\mathcal{K}(X_{0})}\mathcal{J}_{X_{0}}([\omega])=4\pi.$
Subcase 1b: $g=0$. Now
$\phi_{e,0}(t)=\begin{cases}(t-e)\dfrac{2t+2-e}{2t-e},&e<t\leq e+1,\\[6.0pt] \dfrac{2t+2-e}{2t-e},&t\geq e+1.\end{cases}$
For $t\geq e+1$,
$\phi_{e,0}^{\prime}(t)=\frac{d}{dt}\left(\frac{2t+2-e}{2t-e}\right)=-\frac{4}{(2t-e)^{2}}<0,$
so $\phi_{e,0}$ is strictly decreasing on $e+1,+\infty)$, and
$\max_{t\geq e+1}\phi_{e,0}(t)=\phi_{e,0}(e+1)=\frac{e+4}{e+2}.$
For $e<t\leq e+1$, a direct computation gives $\phi_{e,0}^{\prime}(t)>0$, so $\phi_{e,0}$ is strictly increasing on $(e,e+1]$, and
$\max_{e<t\leq e+1}\phi_{e,0}(t)=\phi_{e,0}(e+1)=\frac{e+4}{e+2}.$
Thus
$\sup_{t>e}\phi_{e,0}(t)=\phi_{e,0}(e+1)=\frac{e+4}{e+2},$
and the supremum is attained precisely at $t=e+1$, i.e. on classes proportional to $C_{0}+(e+1)F$.
Case 2: $g=1$ and e=1e=-1. This is the exceptional indecomposable elliptic ruled surface. In this case the effective cone is generated by the classes $F,\penalty 10000\ C_{0},$ and $2C_{0}-F=-K_{X_{0}}$. Hence for a class $\omega]=aC_{0}+bF$, we have
$\omega]\cdot F=a,\qquad[\omega]\cdot C_{0}=a+b,\qquad[\omega]\cdot(2C_{0}-F)=a+2b.$
Then the Kähler condition becomes a>0,b>a/2a>0,\ b>-a/2. Writing $t=b/a$, with $t>-\frac{1}{2}$, we obtain
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=a\min\{1,\ 1+2t\}.$
Moreover, $c_{1}(X_{0})=2C_{0}-F$, so
$c_{1}(X_{0})\cdot[\omega]=2b+a,\qquad[\omega]^{2}=a(2b+a).$
Therefore
$\mathcal{J}_{X_{0}}([\omega])=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot 4\pi\frac{2b+a}{a(2b+a)}=4\pi\min\{1,\ 1+2t\}.$
It follows immediately that
$\sup_{\mathcal{K}(X_{0})}\mathcal{J}_{X_{0}}([\omega])=4\pi,$
and the supremum is attained precisely when $t\geq 0$ that is, $\frac{b}{a}\geq 0$. Combining the above cases proves the proposition. ∎
Remark 4.3.
When $B\cong\mathbb{P}^{1}$, the ruled surface $X_{0}$ is the Hirzebruch surface $\mathbb{F}_{e}$ with $e\geq 0$. In particular, if $e=1$, $\mathbb{F}_{1}$ is not minimal and is isomorphic to the blowup $\operatorname{Bl}_{1}\mathbb{P}^{2}$ of $\mathbb{P}^{2}$ at one point. As expected, our computations of $\mathcal{J}_{\operatorname{Bl}_{1}\mathbb{P}^{2}}$ in Example 3.5 and of $\mathcal{J}_{\mathbb{F}_{1}}$ in Proposition 4.2 yield the same value..
For later use, we also fix notation for blow-ups of ruled surfaces. Let $X_{k}=\operatorname{Bl}_{k}X_{0}$ be the blow-up of $X_{0}$ at $k$ points in very general position, and denote by $E_{1},\dots,E_{k}$ the exceptional divisors. Here “very general position” means that the blown-up points lie on distinct fibres of the ruling, and none of them lies on the distinguished section $C_{0}$. This is the worst configuration for our purposes: if some points lie on the same fibre, or on $C_{0}$, additional effective curves appear whose strict transforms can only decrease $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$, and hence can only decrease $\mathcal{J}_{X_{k}}([\omega])$.
On $X_{k}$, $k\geq 1$, we have
$NS^{1}(X_{k};\mathbb{R})=\langle C_{0},F,E_{1},\dots,E_{k}\rangle,$
with intersection pairings
$C_{0}^{2}=-e,\qquad F^{2}=0,\qquad C_{0}\cdot F=1,$
and
$E_{i}^{2}=-1,\qquad E_{i}\cdot C_{0}=E_{i}\cdot F=E_{i}\cdot E_{j}=0\quad(i\neq j).$
We will write a Kähler class on $X_{k}$ in the form
$\omega]=aC_{0}+bF-\sum_{i=1}^{k}t_{i}E_{i},\qquad a>0,\quad t_{i}>0.$
The first Chern class is
$c_{1}(X_{k})=\pi^{*}c_{1}(X_{0})-\sum_{i=1}^{k}E_{i}=2C_{0}+(2-2g+e)F-\sum_{i=1}^{k}E_{i}.$
Consequently,
$\omega]^{2}=2ab-ea^{2}-\sum_{i=1}^{k}t_{i}^{2},$
and
$c_{1}(X_{k})\cdot[\omega]=2b+(2-2g-e)a-\sum_{i=1}^{k}t_{i}.$
These formulas will be used repeatedly in the subsequent subsections.
4.2. Non-rational ruled surfaces
Suppose $X_{0}\to B$ is a compact ruled surface over a smooth complex curve $B$ of genus $g>0$. It follows from Proposition 4.2 that
$\sup_{\mathcal{K}(X_{0})}\mathcal{J}_{X_{0}}([\omega])=4\pi.$
We now study blow-ups of such ruled surfaces.
Lemma 4.4.
Let $X_{0}$ be a minimal ruled surface over a smooth curve $B$ of genus $g\geq 1$, with invariant $e\geq-1$. For an integer $k\geq 1$, let $X_{k}=\text{Bl}_{k}(X_{0})$ be the blow-up at $k$ points in very general position, with exceptional divisors $E_{1},\dots,E_{k}$. Then for every $k\geq 1$ one has
$\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}([\omega])=2\pi.$
Proof.
Since $\mathcal{J}_{X_{k}}([\omega])$ is homogeneous of degree 0 in $\omega]$, we may normalize $a=1$ and write
$\omega]=C_{0}+bF-\sum_{i=1}^{k}t_{i}E_{i},\qquad t_{i}>0,$
and $b>e$ if $e\geq 0$; b>1/2b>-1/2 if $g=1$ and e=1e=-1. In particular, since $\omega]$ is Kähler, we always have $\omega]^{2}=2b-e-\sum t_{i}^{2}>0$.
As in the minimal case, $\mathcal{J}_{X_{k}}([\omega])$ can be written as
$\mathcal{J}_{X_{k}}([\omega])=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot 4\pi\,\frac{2b+2-2g-e-\sum_{i=1}^{k}t_{i}}{2b-e-\sum_{i=1}^{k}t_{i}^{2}}.$
Since all blowup points are in very general positions, there is no $p_{i}$ lies on $C_{0}$ and no two $p_{i}$ lie on the same fibre. For each blownup point $p_{i}\in X_{0}$, let $F_{i}$ be the unique fibre through $p_{i}$ and let $\widetilde{F}_{i}\subset X_{k}$ be its strict transform. Then $\widetilde{F}_{i}$ has class $F-E_{i}$ and is effective. Since each exceptional divisor $E_{i}$ is also effective, we obtain
$\omega]\cdot E_{i}=t_{i},\qquad[\omega]\cdot(F-E_{i})=(C_{0}+bF-\sum t_{j}E_{j})\cdot(F-E_{i})=1-t_{i}.$
In particular,
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\leq\min_{1\leq i\leq k}\bigl\{[\omega]\cdot E_{i},\ [\omega]\cdot(F-E_{i})\bigr\}=\min_{1\leq i\leq k}\{t_{i},1-t_{i}\}.$
The Kähler condition $\omega]\cdot(F-E_{i})>0$ implies $0<t_{i}<1$, hence for each $i$ the function $\min\{t_{i},1-t_{i}\}$, viewed as a function of $t_{i}\in(0,1)$, attains its maximal value $1/2$ at $t_{i}=1/2$. Thus for all Kähler classes
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\leq\frac{1}{2}.$
On the other hand, using $g\geq 1$ and $0<t_{i}<1$ we have
$\bigl(2b+2-2g-e-\sum t_{i}\bigr)-\bigl(2b-e-\sum t_{i}^{2}\bigr)=(2-2g)+\sum(t_{i}^{2}-t_{i})<0.$
Hence
$\frac{2b+2-2g-e-\sum t_{i}}{2b-e-\sum t_{i}^{2}}<1,$
and therefore
$\mathcal{J}_{X_{k}}([\omega])\leq 4\pi\,\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\leq 4\pi\cdot\frac{1}{2}=2\pi.$
It remains to show that the bound $2\pi$ is sharp. Fix $t_{1}=\cdots=t_{k}=1/2$ and consider the family of classes
$\omega_{b}]=C_{0}+bF-\frac{1}{2}\sum_{i=1}^{k}E_{i},\qquad b\gg 1.$
For $b$ sufficiently large, these classes lie in the Kähler cone.
We claim that, for all sufficiently large $b$,
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{b}])=\frac{1}{2}.$
Indeed, each exceptional divisor $E_{i}$ is effective and satisfies $\omega_{b}]\cdot E_{i}=1/2$. Moreover, since the blown-up points lie on distinct fibres, the strict transform of the fibre through $p_{i}$ is effective of class $F-E_{i}$, and $\omega_{b}]\cdot(F-E_{i})=1/2.$ Hence
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{b}])\leq\frac{1}{2}.$
Conversely, let $C\subset X_{k}$ be any irreducible effective curve distinct from the $E_{i}$ and the strict transforms $F-E_{i}$. Write its numerical class as
$C]=dC_{0}+nF-\sum_{i=1}^{k}m_{i}E_{i},\qquad d,n,m_{i}\in\mathbb{Z}_{\geq 0}.$
Then
$\omega_{b}]\cdot C=(b-e)d+n-\frac{1}{2}\sum_{i=1}^{k}m_{i}.$
If $d\geq 1$, this tends to $\infty$ as $b\to+\infty$, regardless of whether $e\geq 0$ or e=1e=-1. Thus it is strictly bigger than $1/2$ for all sufficiently large $b$. If $d=0$, then $C$ is supported in fibres and exceptional divisors. Since the blown-up points lie on distinct fibres, the only irreducible possibilities are $E_{i}$ and $F-E_{i}$, which have already been treated. Therefore every other irreducible effective curve has intersection strictly bigger than $1/2$ for $b\gg 1$, and hence
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{b}])=\frac{1}{2}.$
For these classes we have
$\mathcal{J}_{X_{k}}([\omega_{b}])=\frac{1}{2}\cdot 4\pi\,\frac{2b+2-2g-e-k/2}{2b-e-k/4}.$
As $b\to+\infty$, the scalar-curvature factor tends to $4\pi$, and therefore
$\mathcal{J}_{X_{k}}([\omega_{b}])\nearrow 2\pi.$
This proves
$\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}([\omega])=2\pi.$
Combining the estimates for the minimal ruled surface $X_{0}$ and its blowups $X_{k}$ we obtain the following systolic inequality in the non-rational case.
Theorem 4.5.
Let $X\to B$ be a ruled surface (not necessarily minimal) fibred over a complex curve $B$ of genus $g\geq 1$, and let $\omega$ be a PSC Kähler metric on $X$. Then
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)\leq 4\pi.$
Moreover, equality holds if and only if $B$ is an elliptic curve, $X$ is minimal and isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$, so that the induced metric $\omega$ has constant scalar curvature with holomorphic $2$-systole realized by the $\mathbb{P}^{1}$-fibre.
Proof.
By Theorem 2.2 (1), Proposition 4.2, and Lemma 4.4, every PSC Kähler metric $\omega$ on a ruled surface over a base of genus $g\geq 1$ satisfies
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)\leq\mathcal{J}_{X}([\omega])\leq 4\pi.$
Moreover, if $X$ is non-minimal, then Lemma 4.4 gives the sharper bound
$\mathcal{J}_{X}([\omega])\leq 2\pi,$
so equality in
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)\leq 4\pi$
can occur only when $X$ is minimal. By Proposition 4.2, the minimal case attains the value $4\pi$ only when $g=1$. Equality in Theorem 2.2 (1) then forces $\omega$ to have constant scalar curvature and the holomorphic $2$-systole to be realized by the $\mathbb{P}^{1}$-fibre.
It remains to prove that, in this equality case, $X$ is isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$. Since $X\xrightarrow{\pi}B$ is a minimal ruled surface over an elliptic curve, there exists a rank-$2$ holomorphic vector bundle $\mathcal{E}$ on $B$ such that $X=\mathbb{P}(\mathcal{E})$ and $\pi$ is the bundle projection (see [14, Chapter V]). By the classification of cscK metrics on ruled surfaces due to Apostolov and Tønnesen-Friedman [1, Theorem 2], the existence of a cscK metric on $X=\mathbb{P}(\mathcal{E})$ is equivalent to the slope-polystability of $\mathcal{E}$. Hence $\mathcal{E}$ is polystable. By Fujiki’s equivalences [10, Lemma 2], this implies that $X$ is quasi-stable. Equivalently, there exists a projective unitary representation
$\rho:\pi_{1}(B)\longrightarrow PU(2)\subset PGL_{2}(\mathbb{C})$
such that $X$ is biholomorphic to the suspension quotient
$X_{\rho}:=(\widetilde{B}\times\mathbb{P}^{1})/\pi_{1}(B),\qquad\gamma\cdot(z,[v])=(\gamma z,\rho(\gamma)[v]),$
where $\widetilde{B}\cong\mathbb{C}$ is the universal cover. Fix a biholomorphism $\Phi:X\stackrel{{\scriptstyle\sim}}{{\to}}X_{\rho}$, and set
$\omega_{\rho}:=(\Phi^{-1})^{*}\omega.$
Then $\omega_{\rho}$ is a cscK Kähler metric on $X_{\rho}$, and $\Phi$ is an isometry between $(X,\omega)$ and $(X_{\rho},\omega_{\rho})$.
Since $g(B)=1$, Fujiki’s rigidity result [10, Lemma 10] shows that any cscK Kähler metric on $X_{\rho}$ is a generalized Kähler–Einstein metric in the sense of [10, §3]. It follows that $X_{\rho}$ is isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$ endowed with the product Kähler metric $\omega_{\mathrm{FS}}\oplus\omega_{\mathrm{flat}}$. Therefore $X$ is also isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$. ∎
4.3. Rational ruled surfaces
In this subsection we treat the case where the base curve is rational, so that $B\cong\mathbb{P}^{1}$ and $X_{0}=\mathbb{F}_{e}\longrightarrow\mathbb{P}^{1}$ is a Hirzebruch surface with invariant $e\geq 0$. We denote by $X_{k}=\text{Bl}_{k}(\mathbb{F}_{e})$ the blow-up of $\mathbb{F}_{e}$ at $k$ points in very general positions.
Proposition 4.6.
Let $\mathbb{F}_{e}$ be a Hirzebruch surface with invariant $e\geq 0$, and let $X_{1}=\mathrm{Bl}_{p}(\mathbb{F}_{e})$ be its blow-up at one point. Then:
  1. (1)
    If $e=0$ (so $X_{0}\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$), one has
    $\sup_{\mathcal{K}^{+}(X_{1})}\mathcal{J}_{X_{1}}([\omega])\;=\;4\pi.$
  2. (2)
    If $e\geq 1$, one has
    $\sup_{\mathcal{K}^{+}(X_{1})}\mathcal{J}_{X_{1}}([\omega])\;=\;4\pi\cdot\frac{2e+5}{4e+3}.$
Proof.
We treat separately the cases $e=0$ and $e\geq 1$. When $e=0$, in which case $\mathbb{F}_{0}\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$. Write the two rulings as $F_{1}:=[\mathbb{P}^{1}\times\{\mathrm{pt}\}],\penalty 10000\ F_{2}:=[\{\mathrm{pt}\}\times\mathbb{P}^{1}]$, so that $F_{1}^{2}=F_{2}^{2}=0$, and $F_{1}\cdot F_{2}=1$. Any Kähler class on $X_{1}$ can be written as $\omega=aF_{1}+bF_{2}-tE$. Note that the surface $X_{1}$ is the degree $7$ del Pezzo surface and its Mori cone $\overline{\mathrm{NE}}(X_{1})$ is generated by the three (1)(-1)-curves $E,\penalty 10000\ F_{1}-E$, and $F_{2}-E$. Hence,
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=\min\{t,\ a-t,\ b-t\},$
and
$\mathcal{J}_{X_{1}}([\omega])=\min\{t,\ a-t,\ b-t\}\cdot\frac{4\pi(2a+2b-t)}{2ab-t^{2}}.$
Setting $x:=a/t,\penalty 10000\ y:=b/t$, so that $a=xt$, $b=yt$, then the Kähler conditions $t>0$, at>0a-t>0, bt>0b-t>0, and $\omega]^{2}>0$ translate into $x>1,\quad y>1,$ and $2xy>1$. Thus,
$\mathcal{J}_{X_{1}}([\omega])=4\pi\,\min\{1,x-1,y-1\}\frac{2x+2y-1}{2xy-1},$
defined on $\{(x,y)\mid x>1,\ y>1\}$. A straightforward piecewise analysis of the two-variable function
$(x,y)\longmapsto\min\{1,x-1,y-1\}\frac{2x+2y-1}{2xy-1},\qquad x>1,\ y>1,$
shows that its supremum is $1$, attained uniquely at $x=y=2$. We then conclude
$\sup_{[\omega]}\mathcal{J}_{X_{1}}([\omega])=4\pi\sup_{x>1,y>1}\left(\min\{1,x-1,y-1\}\frac{2x+2y-1}{2xy-1}\right)=4\pi.$
Moreover, the supremum is achieved by the class $\omega]=F_{1}+F_{2}-\tfrac{1}{2}E$. This proves (1).
We now assume $e\geq 1$ and write the Kähler class on $X_{1}$ as $\omega]=aC_{0}+bF-tE,\penalty 10000\ a>0,\ b>0,\ t>0$. In this case, he Mori cone $\overline{\mathrm{NE}}(X_{1})$ is generated by three extremal rays $E$, FEF-E and $C_{0}$. Hence,
$\operatorname{sys}_{2}([\omega])=\min\{t,\ a-t,\ b-ea\}.$
Introduce the scale-invariant variables again $x:=a/t$, and $y:=b/t$, so that $a=xt$, $b=yt$. The Kähler inequalities $t>0$, at>0a-t>0, bea>0b-ea>0 are equivalent to
$x>1,\qquad y>ex.$
It is convenient to write $y=ex+z$ with $z>0$, so that
$\mathcal{J}_{X_{1}}([\omega])=4\pi\,\min\{1,\ x-1,\ z\}\frac{(e+2)x+2z-1}{ex^{2}+2xz-1}:=4\pi\penalty 10000\ \phi_{e}(x,z),$
defined on the domain
$\mathcal{D}_{e}:=\{(x,z)\in\mathbb{R}^{2}\mid x>1,\ z>0,\ ex^{2}+2xz-1>0\}.$
The function $\phi_{e}(x,z)$ is piecewise rational, according to the regions where $\min\{1,x-1,z\}$ is realized by one of its three entries. A straightforward inspection on these regions shows that $\phi_{e}$ attains its global maximum at $(x,z)=(2,1)$, with value
$\phi_{e}(2,1)=\frac{2e+5}{4e+3}.$
The corresponding ray in the Kähler cone is given by $\omega]=2t\,C_{0}+(2e+1)t\,F-tE$, and we obtain
$\sup_{\mathcal{K}^{+}(X_{1})}\mathcal{J}_{X_{1}}([\omega])=4\pi\cdot\frac{2e+5}{4e+3}.$
This proves (2) and completes the proof of the lemma. ∎
Having dealt with the case of a single blow-up of a Hirzebruch surface, we now turn to blowing up several points on $\mathbb{F}_{0}\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$. As in the $\mathbb{P}^{2}$-case (see Proposition 3.1 in Section 3), the key input is a simple “mass-shift” optimization for the exceptional parameters $t_{i}$, which allows us to reduce to an extremal configuration and then perform a purely algebraic estimate.
Lemma 4.7.
Let $X_{0}=\mathbb{P}^{1}\times\mathbb{P}^{1}$ and let $X_{k}=\text{Bl}_{k}(X_{0})$ be the blow-up of $X_{0}$ at $k\geq 2$ points in very general position. Then
$\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}([\omega])\;\leq\;4\pi.$
Proof.
Let $F_{1}=[\mathbb{P}^{1}\times\{\mathrm{pt}\}]$ and $F_{2}=[\{\mathrm{pt}\}\times\mathbb{P}^{1}]$ be the two rulings, and let $E_{1},\dots,E_{k}$ be the exceptional curves. By the symmetry between the two rulings and the scale invariance of $\mathcal{J}_{X_{k}}$, we may, after possibly interchanging $F_{1}$ and $F_{2}$ and rescaling, assume that the coefficient of the smaller ruling is $1$. More precisely, any Kähler class can be written (after possibly interchanging $F_{1},F_{2}$ and rescaling) in the form $\omega]=F_{1}+bF_{2}-\sum t_{i}E_{i},$ for $b\geq 1,\ t_{i}>0$. Set $m=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$. The effective curves $E_{i},\ F_{1}-E_{i},\ F_{2}-E_{i}$ give candidates for $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$, hence
$\omega]\cdot E_{i}=t_{i}\geq m,\quad[\omega]\cdot(F_{1}-E_{i})=1-t_{i}\geq m,\quad[\omega]\cdot(F_{2}-E_{i})=b-t_{i}\geq m.$
Hence, for all $i$,
$m\;\leq\;t_{i}\;\leq\;1-m,\qquad t_{i}\;\leq\;b-m.$ (4.2)
In particular $0<m\leq\tfrac{1}{2}$ and $b\geq 1$.
In this setting, we again apply the mass–shifting reduction introduced in the $\mathbb{P}^{2}$ case. We will not elaborate further on this point and refer the reader to Section 3 for details. Write $S:=\sum_{i=1}^{k}t_{i}$ and $Q:=\sum_{i=1}^{k}t_{i}^{2}$. Thus
$\mathcal{J}_{X_{k}}([\omega])=4\pi\,m\cdot\frac{2b+2-S}{2b-Q}.$ (4.3)
Fix $(b,m,S)$ in the range $0<m\leq\tfrac{1}{2},\quad b\geq 1,\quad km\leq S\leq 1+(k-2)m$. For such parameters any admissible $k$-tuple $t=(t_{1},\dots,t_{k})$ satisfies $t_{i}\geq m$ and $\sum t_{i}=S$. By Proposition 3.1, under the constraints $t_{i}\geq m,\ \sum t_{i}=S$, the quadratic form $Q(t)=\sum t_{i}^{2}$ is maximized precisely when
$t=(\underbrace{m,\dots,m}_{k-1},\,S-(k-1)m)$
up to permutation, and in that case
$Q_{\max}=(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2}.$
Consequently, for each fixed triple $(b,m,S)$, we have
$\mathcal{J}_{X_{k}}([\omega])\;\leq\;4\pi\,m\cdot\frac{2b+2-S}{2b-Q_{\max}}=:4\pi\,\phi(b,m,S),$
where
$\phi(b,m,S)=\frac{m(2b+2-S)}{2b-(k-1)m^{2}-\bigl(S-(k-1)m\bigr)^{2}}.$ (4.4)
Set
$N:=m(2b+2-S),\qquad D:=2b-(k-1)m^{2}-\bigl(S-(k-1)m\bigr)^{2},$
so that $\phi=N/D$ on the domain $\{D>0\}\cap\{N>0\}$. We claim that
$D(b,m,S)\;\geq\;N(b,m,S)$
for all admissible $(b,m,S)$, which immediately implies $\phi(b,m,S)\leq 1$.
To verify the claim, a direct computation gives
$D-N$ $=2b-(k-1)m^{2}-\bigl(S-(k-1)m\bigr)^{2}-m(2b+2-S)$
$=2b(1-m)-2m+m(S-(k-1)m)-\left(S-(k-1)m\right)^{2}.$
Introduce z:=S(k1)mz:=S-(k-1)m with $m\leq z\leq 1-m$. In terms of $(m,z)$ we obtain
$D-N=2b(1-m)-2m+mz-z^{2}.$
Observe that DND-N is strictly increasing as a function of $b$. Under the constraint $b\geq 1$ it attains its minimum at $b=1$. Thus
$D(b,m,S)-N(b,m,S)\;\geq\;G(m,z),$
where
$G(m,z):=D(1,m,S)-N(1,m,S)=2-4m+mz-z^{2}.$
On the interval $z\in[m,1-m]$ we have $\partial_{z}G(m,z)=m-2z<0$, so $G(m,\cdot)$ is strictly decreasing, and hence
$\min_{z\in[m,1-m]}G(m,z)=G(m,1-m).$
A short calculation yields
$G(m,1-m)=-(2m^{2}+m-1)\geq 0.$
Therefore $G(m,z)\geq 0$ for all $z\in[m,1-m]$. Combining this with the monotonicity in $b$ gives
$D(b,m,S)-N(b,m,S)\;\geq\;0$
for every admissible triple $(b,m,S)$, and hence $\phi(b,m,S)\leq 1$.
Consequently,
$\mathcal{J}_{X_{k}}([\omega])\;\leq\;4\pi\,\phi(b,m,S)\;\leq\;4\pi$
for Kähler class $\omega]\in\mathcal{K}^{+}(X_{k})$ on $X_{k}$. This shows that
$\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}([\omega])\leq 4\pi,$
which is the desired estimate. ∎
We now consider blow-ups of the general Hirzebruch surfaces $\mathbb{F}_{e}$ with $e\geq 1$. The argument is parallel, with the only new feature being the contribution of the negative section $C_{0}$ in the intersection computations. The final bound is again $4\pi$, and when $e=1$, this recovers the case of $\mathbb{P}^{2}$ blown up at one point via the identification $\mathbb{F}_{1}\cong\mathbb{P}^{2}\#\overline{\mathbb{P}^{2}}$.
Lemma 4.8.
Let $X_{0}=\mathbb{F}_{e}$ be a Hirzebruch surface with $e\geq 1$, and let $X_{k}=\text{Bl}_{k}(X_{0})$ be the blow-up of $X_{0}$ at $k\geq 2$ points in very general position. Then
$\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}([\omega])\;\leq\;4\pi.$
Proof.
After rescaling, write any Kähler class on $X_{k}$ in the form $\omega]\;=\;C_{0}+bF-\sum t_{i}E_{i}$, for $b>e,\;t_{i}>0$. We use the same notations as before, so that
$\mathcal{J}_{X_{k}}([\omega])=4\pi\,m\cdot\frac{2b+2-e-S}{2b-e-Q}.$ (4.5)
Fix parameters $(b,m,S)$ in the range
$0<m\leq\tfrac{1}{2},\quad b\geq e+m,\quad km\leq S\leq 1+(k-2)m.$
By the mass–shift Proposition 3.1, under the constraints $t_{i}\geq m,\ \sum t_{i}=S$, the quadratic form $Q(t)=\sum t_{i}^{2}$ is maximized precisely when
$t=(\underbrace{m,\dots,m}_{k-1},\,S-(k-1)m)$
up to permutation, and in that case
$Q_{\max}=(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2}.$
Hence, for each fixed triple $(b,m,S)$,
$\mathcal{J}_{X_{k}}([\omega])\;\leq\;4\pi\,m\cdot\frac{2b+2-e-S}{2b-e-Q_{\max}}=:4\pi\,\phi_{e,k}(b,m,S),$
where
$\phi_{e,k}(b,m,S):=\frac{m\bigl(2b+2-e-S\bigr)}{2b-e-(k-1)m^{2}-\bigl(S-(k-1)m\bigr)^{2}}.$ (4.6)
We use the same trick as in the proof of Lemma 4.7 with just a slight difference when we deal with the extremum. Set
$N:=m\bigl(2b+2-e-S\bigr),\qquad D:=2b-e-(k-1)m^{2}-\bigl(S-(k-1)m\bigr)^{2},$
so that $\phi_{e,k}=N/D$ on the domain $\{D>0\}\cap\{N>0\}$. We claim that
$D(b,m,S)\;\geq\;N(b,m,S)$
for all admissible $(b,m,S)$, which immediately implies $\phi_{e,k}(b,m,S)\leq 1$.
A direct computation gives
$D-N$ $=2b-e-(k-1)m^{2}-\bigl(S-(k-1)m\bigr)^{2}-m\bigl(2b+2-e-S\bigr)$ (4.7)
$=2b(1-m)-e(1-m)-(k-1)m^{2}-\bigl(S-(k-1)m\bigr)^{2}-2m+mS.$
Note that DND-N is strictly increasing as a function of $b$. Under the constraint $b\geq e+m$ it attains its minimum at $b_{0}:=e+m$. In particular,
$D(b,m,S)-N(b,m,S)\;\geq\;D(b_{0},m,S)-N(b_{0},m,S).$ (4.8)
Substituting $b=b_{0}=e+m$ into (4.7) yields
$D(b_{0},m,S)-N(b_{0},m,S)=e(1-m)-2m^{2}+mS-\bigl(S-(k-1)m\bigr)^{2}.$
Introduce z:=S(k1)mz:=S-(k-1)m with $m\ \leq\ z\ \leq\ 1-m$. In terms of $(m,z)$ we obtain
$D(b_{0},m,S)-N(b_{0},m,S)=F_{e}(m,z):=e(1-m)-2m^{2}+mz-z^{2}.$ (4.9)
Combining (4.8) and (4.9) we deduce
$D(b,m,S)-N(b,m,S)\;\geq\;F_{e}(m,z)$
for all admissible $(b,m,S)$. A short computation gives
$\min_{z\in[m,1-m]}F_{e}(m,z)=F_{e}(m,1-m)$ $=e(1-m)-2m^{2}+m(1-m)-(1-m)^{2}$
$=e(1-m)-4m^{2}+3m-1$
$=(e-1)+(3-e)m-4m^{2}$
$\geq 0,$
with strict inequality if $e\geq 2$. In particular, $F_{e}(m,z)\geq 0$ for all admissible $m,z$, and hence $\phi_{e,k}(b,m,S)\leq 1$.
Consequently,
$\mathcal{J}_{X_{k}}([\omega])\;\leq\;4\pi\,\phi_{e,k}(b,m,S)\;\leq\;4\pi$
for every Kähler class $\omega]$ on $X_{k}$. Therefore
$\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}([\omega])\leq 4\pi,$
as claimed. ∎
Before turning to the global statement for rational ruled surfaces, it is instructive to isolate a concrete situation where the quantity $\mathcal{J}_{X_{k}}$ can be written down explicitly. In the regime of blowing up at most $e$ points on a Hirzebruch surface (namely $k\leq e$) in very general position, the Mori cone is finitely generated by a short list of curves, so that $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$ and even $\mathcal{J}_{X_{k}}([\omega])$ reduce to an explicit finite-dimensional optimization problem in the parameters of the Kähler class. The following example makes this reduction precise.
Example 4.9.
Fix an integer $e\geq 1$ and let $X_{0}=\mathbb{F}_{e}$ be the $e$-th Hirzebruch surface with section $C_{0}$ and fibre class $F$. Let $X_{k}:=\text{Bl}_{k}(\mathbb{F}_{e})$ be the blow-up of $k$ points $p_{1},\dots,p_{k}$ with $k\leq e$ in very general position. In this setting, the Mori cone $\overline{\mathrm{NE}}(X_{k})$ is polyhedral. Moreover, see for instance [13, Proposition 2.4 and Lemma 3.4], every extremal ray of $\overline{\mathrm{NE}}(X_{k})$ is generated by $C_{0},\penalty 10000\ E_{i},\penalty 10000\ F-E_{i}$ for $1\leq i\leq k$. We now fix a Kähler class on $X_{k}$ and normalize it as in Lemma 4.8:
$\omega]=C_{0}+bF-\sum_{i=1}^{k}t_{i}E_{i},\qquad b>e,\ t_{i}>0.$
Using the intersection form on $X_{k}$ we obtain
$m:=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=\min\bigl\{t_{i},\ 1-t_{i},\ b-e:\ 1\leq i\leq k\bigr\}.$
From $t_{i}\geq m$ and $1-t_{i}\geq m$ we immediately deduce $0<m\leq\frac{1}{2}$. The functional $\mathcal{J}_{X_{k}}([\omega])$ takes the explicit form
$\mathcal{J}_{X_{k}}([\omega])=4\pi\,m\cdot\frac{2b+2-e-S}{2b-e-Q}.$
and the parameters satisfy
$0<m\leq\tfrac{1}{2},\quad 0<t_{i}<1,\quad b>e,\quad 2b-e-Q>0.$
Assume in addition that $k\leq 8$. Then a direct optimization of the above function under these constraints shows that the maximum is achieved at (e+1/2,1/2,,1/2)(e+1/2,1/2,...,1/2). Consequently,
$\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}([\omega])=4\pi\,\frac{2e+6-k}{4e+4-k}\leq 4\pi,\qquad k\leq\min\{e,8\}.$
Remark 4.10.
The threshold $k=8$ in Example 4.9 has a genuine structural meaning. From the algebro-geometric point of view, since $K_{\mathbb{F}_{e}}^{2}=8$ and each blow-up decreases $K^{2}$ by $1$, one has $K_{X_{k}}^{2}=8-k$. Thus $k=8$ is precisely the point where the sign of $K_{X_{k}}^{2}$ changes. This is the analogue, for blow-ups of Hirzebruch surfaces, of the classical transition at $k=9$ for blow-ups of $\mathbb{P}^{2}$.
The same threshold also appears directly in the optimization of $\mathcal{J}_{X_{k}}$. Indeed, writing
$\Phi(b,t):=m\,\frac{2b+2-e-S}{2b-e-Q},\qquad m=\min\{t_{i},1-t_{i},b-e\},$
one computes
$\partial_{b}\Phi=\frac{2m(S-Q-2)}{(2b-e-Q)^{2}},\qquad S-Q=\sum_{i}t_{i}(1-t_{i})\leq\frac{k}{4}.$
[…]
Combining the previous lemmas, we obtain the following theorem.
Theorem 4.11.
Let $X\to\mathbb{P}^{1}$ be a rational ruled surface (not necessarily minimal) endowed with a PSC Kähler metric $\omega$. Then
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)\leq 8\pi.$
Moreover, equality holds if and only if $X\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$, endowed with the product Fubini-Study metric.
Proof.
Let $X\to\mathbb{P}^{1}$ be a rational ruled surface endowed with a PSC Kähler metric $\omega$. Since every rational ruled surface is obtained from a Hirzebruch surface $\mathbb{F}_{e}$ by blowing up finitely many points, we distinguish cases according to the number of blow-ups.
If $X$ is minimal, then $X\cong\mathbb{F}_{e}$ for some $e\geq 0$. By Proposition 4.2,
$\sup_{\mathcal{K}(\mathbb{F}_{e})}\mathcal{J}_{\mathbb{F}_{e}}([\omega])=4\pi\,\frac{e+4}{e+2}\leq 8\pi,$
with equality if and only if $e=0$, i.e. $X\cong\mathbb{F}_{0}\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$.
If $X$ is the blow-up of $\mathbb{F}_{e}$ at one point, then Proposition 4.6 gives
$\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])=\begin{cases}4\pi,&e=0,\\[4.0pt] 4\pi\,\dfrac{2e+5}{4e+3},&e\geq 1,\end{cases}$
and in either case this is strictly smaller than $8\pi$.
If $X$ is the blow-up of $\mathbb{F}_{e}$ at $k\geq 2$ points, then Lemma 4.7 and 4.8 gives
$\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])\leq 4\pi.$
Combining all cases, we obtain
$\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])\leq 8\pi,$
and equality can occur only when $X\cong\mathbb{F}_{0}\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$. Now apply Theorem 2.2 (1). Since $\omega$ is PSC, we have $\omega]\in\mathcal{K}^{+}(X)$, and therefore
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)\leq\mathcal{J}_{X}([\omega])\leq 8\pi.$
This proves the inequality.
It remains to characterize the equality case. Suppose $\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)=8\pi$. Then necessarily $\mathcal{J}_{X}([\omega])=8\pi$, so by the discussion above we must have $X\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$. Moreover, equality in Theorem 2.2 (1) implies that $\omega$ is cscK and that $\operatorname{sys}_{2}(\omega)$ is realized by a holomorphic curve. On $\mathbb{P}^{1}\times\mathbb{P}^{1}$, write $\omega]=aF_{1}+bF_{2}$, for $a,b>0$ where $F_{1},F_{2}$ are the two rulings. Assuming without loss of generality that $a\leq b$, then
$\mathcal{J}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}([\omega])=4\pi\,a\,\frac{a+b}{ab}=4\pi\left(1+\frac{a}{b}\right)\leq 8\pi,$
with equality if and only if $a=b$. Thus $\omega]$ is proportional to $F_{1}+F_{2}$. Since $\omega$ is cscK, it follows that, up to overall scaling, $\omega$ is the product Fubini–Study metric on $\mathbb{P}^{1}\times\mathbb{P}^{1}$.
Conversely, the product Fubini-Study metric on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ has constant scalar curvature, and its $2$-systole is realized by either ruling $\mathbb{P}^{1}$. A direct computation shows that equality holds:
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)=8\pi.$
This completes the proof. ∎
5. Level set method on non-rational PSC Kähler surfaces
In [22], Stern introduced the following inequality for a non-constant $S^{1}$-valued harmonic map $u$ on a $3$-manifold $(M,g)$ through the level set method:
$2\pi\int_{\theta\in S^{1}}\chi\left(\Sigma_{\theta}\right)d\theta\;\geq\;\frac{1}{2}\int_{\theta\in S^{1}}\left[\int_{\Sigma_{\theta}}\left(|du|^{-2}|\operatorname{Hess}(u)|^{2}+\text{scal}_{M}(g)\right)dV_{g}\right]d\theta,$ (5.1)
and used it to give a new proof of the Bray–Brendle–Neves systolic inequality for the $2$-systole. In this section, we adapt the level set method to non-rational PSC Kähler surfaces, and obtain an alternative proof of Theorem 4.5.
5.1. A Stern-type scalar curvature inequality
Let $X$ be a compact non-rational PSC Kähler surface. By the classification result recalled in the introduction, $X$ is a ruled surface fibred over a compact Riemann surface $B$ of genus $g(B)\geq 1$. Denote by
$f\colon(X,\omega)\longrightarrow B$
the induced non-constant holomorphic fibration. By the uniformization theorem, we may equip $B$ with a constant curvature metric $\omega_{0}$ of non-positive Gaussian curvature (so $\omega_{0}$ is hyperbolic if $g(B)\geq 2$ and flat if $g(B)=1$).
Fix a point $z\in B$. Then $f^{-1}(z)$ is a (possibly singular) Cartier divisor on $X$, which we denote by $D_{z}$. It defines a line bundle $\mathcal{O}(D_{z})$ whose first Chern class is represented by $f^{*}\omega_{0}$ (after a suitable normalization). In what follows we restrict attention to regular values of $f$, so that $D_{z}$ is smooth, and we keep the notation $D_{z}$ for the smooth fibre. Recall that for a smooth divisor $D\subset X$, the adjunction formula states
$K_{D}=\bigl(K_{X}\otimes\mathcal{O}(D)\bigr)|_{D}.$
Since $D=D_{z}$ is a fibre of the holomorphic fibration $f\colon X\to B$, its normal bundle $\mathcal{N}_{D}\cong\mathcal{O}(D)|_{D}$ is holomorphically trivial. By the adjunction formula,
$K_{D}=\bigl(K_{X}\otimes\mathcal{O}(D)\bigr)|_{D}.$
Taking first Chern classes, we obtain
$c_{1}(D)=c_{1}(X)|_{D}.$
Therefore, for the Ricci forms induced by $\omega$, the difference
$\operatorname{Ric}_{D}(\omega)-\operatorname{Ric}_{X}(\omega)|_{D}$
is an exact real $(1,1)$-form on the compact Riemann surface $D$. By the $\partial\bar{\partial}$-lemma, there exists a smooth real-valued function $\phi\in C^{\infty}(D)$ such that
$\operatorname{Ric}_{D}(\omega)=\operatorname{Ric}_{X}(\omega)|_{D}+\sqrt{-1}\,\partial_{D}\bar{\partial}_{D}\phi.$
Tracing with respect to the induced metric $\omega|_{D}$, we get
$S_{D}(\omega)=\operatorname{tr}_{\omega|_{D}}\bigl(\operatorname{Ric}_{X}(\omega)|_{D}\bigr)+\Delta_{D}\phi.$
Let $\nu$ be a local unit normal vector field of type $(1,0)$ along $D$ with respect to $\omega$. Since $D\subset X$ is a complex hypersurface, we have
$\operatorname{tr}_{\omega|_{D}}\bigl(\operatorname{Ric}_{X}(\omega)|_{D}\bigr)=S_{X}(\omega)-\operatorname{Ric}_{X}(\omega)(\nu,\bar{\nu}).$
Hence
$S_{D}(\omega)=S_{X}(\omega)-\operatorname{Ric}_{X}(\omega)(\nu,\bar{\nu})+\Delta_{D}\phi.$
In particular, since
$\nu=\frac{\nabla^{1,0}f}{|\nabla^{1,0}f|},$
we obtain
$\operatorname{Ric}_{X}(\omega)\bigl(\nabla^{1,0}f,\nabla^{0,1}f\bigr)=|\nabla^{1,0}f|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)+\Delta_{D}\phi\bigr).$ (5.2)
We next recall the Bochner formula for holomorphic maps and a co-area formula adapted to the present setting.
Lemma 5.1 (Bochner formula).
[…]
Lemma 5.2 (Co-area formula).
Let $(X^{n},\omega)$ be a compact Kähler manifold and let $(B,\omega_{0})$ be a compact Riemann surface, normalized so that
$\int_{B}\omega_{0}=\frac{1}{n}.$
Let $f\colon X\to B$ be a non-constant holomorphic map, and let $g\in C^{\infty}(X)$. Then
$\int_{X}g\,\omega^{n}=\int_{B}\left(\int_{f^{-1}(z)}\frac{g}{|\partial f|^{2}}\,\omega^{n-1}\right)\omega_{0}.$
When $(N,\tilde{\omega})=(B,\omega_{0})$, the curvature term $\operatorname{tr}^{2}_{\omega}(f^{*}\operatorname{Rm}(\omega_{0}))$ is non-negative and vanishes identically if and only if $(B,\omega_{0})$ is flat. In particular,
$\Delta|\partial f|^{2}\;\geq\;|\nabla\partial f|^{2}+\operatorname{Ric}_{X}(\omega)\bigl(\nabla^{1,0}f,\nabla^{0,1}f\bigr),$
with equality if and only if $(B,\omega_{0})$ is an elliptic curve with a flat metric. Combining this with (5.2) we obtain
$\Delta|\partial f|^{2}\;\geq\;|\nabla\partial f|^{2}+|\nabla^{1,0}f|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)+\Delta_{D}\phi\bigr).$ (5.3)
We now integrate this inequality in the fibre direction using the co-area formula. The following lemma is the basic Stern-type inequality that we shall use in the non-rational ruled case.
Lemma 5.3.
Let $(X^{n},\omega)$ be a compact Kähler manifold and let $(B,\omega_{0})$ be a compact Riemann surface of genus $g(B)\geq 1$, endowed with a constant curvature metric $\omega_{0}$. Suppose that $f\colon X\rightarrow B$ is a non-constant holomorphic map, and let $D=f^{-1}(z)$ denote a regular fibre. Then
$\int_{B}\left[\int_{D}\left(\frac{\left|\nabla\partial f\right|^{2}}{|\partial f|^{2}}+S_{X}(\omega)-S_{D}(\omega)\right)\omega^{n-1}\right]\omega_{0}\;\leq\;0.$ (5.4)
Moreover, equality holds in (5.4) if and only if $g(B)=1$ and $(B,\omega_{0})$ is an elliptic curve with a flat metric.
Proof.
Write $B=A\cup B_{0}$, where $A$ is an open neighbourhood of the finite set of critical values of $f$, and $B_{0}$ consists only of regular values. Integrating (5.3) over $f^{-1}(B_{0})$, we obtain
$\int_{f^{-1}(B_{0})}\Bigl(\bigl|\nabla\partial f\bigr|^{2}+\bigl|\nabla^{1,0}f\bigr|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)+\Delta_{D}\phi\bigr)\Bigr)\omega^{n}\;\leq\;\int_{f^{-1}(B_{0})}\bigl(\Delta|\partial f|^{2}\bigr)\omega^{n}.$ (5.5)
Since $X$ is compact and has no boundary,
$\int_{X}\bigl(\Delta|\partial f|^{2}\bigr)\omega^{n}=0.$
Hence, by choosing $A$ with arbitrarily small measure and using Sard’s theorem, we may pass to the limit in (5.5) and obtain
$\int_{X}\Bigl(\bigl|\nabla\partial f\bigr|^{2}+\bigl|\nabla^{1,0}f\bigr|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)+\Delta_{D}\phi\bigr)\Bigr)\omega^{n}\;\leq\;0.$
Now apply the co-area formula term by term. Since $|\nabla^{1,0}f|^{2}=|\partial f|^{2}$, we get
$\int_{X}|\nabla\partial f|^{2}\,\omega^{n}$ $=n\int_{B}\left(\int_{D}\frac{|\nabla\partial f|^{2}}{|\partial f|^{2}}\,\omega^{n-1}\right)\omega_{0},$
$\int_{X}|\nabla^{1,0}f|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)\bigr)\,\omega^{n}$ $=n\int_{B}\left(\int_{D}\bigl(S_{X}(\omega)-S_{D}(\omega)\bigr)\,\omega^{n-1}\right)\omega_{0},$
and similarly
$\int_{X}|\nabla^{1,0}f|^{2}\,\Delta_{D}\phi\,\omega^{n}=n\int_{B}\left(\int_{D}\Delta_{D}\phi\,\omega^{n-1}\right)\omega_{0}.$
Since each regular fibre $D$ is compact without boundary,
$\int_{D}\Delta_{D}\phi\,\omega^{n-1}=0.$
Therefore
$0$ $\geq n\int_{B}\left[\int_{D}\left(\frac{|\nabla\partial f|^{2}}{|\partial f|^{2}}+S_{X}(\omega)-S_{D}(\omega)\right)\omega^{n-1}\right]\omega_{0},$
which gives (5.4). The equality statement follows from the discussion before (5.3): equality in the Bochner inequality holds if and only if $\operatorname{Rm}(\omega_{0})\equiv 0$, namely if and only if $g(B)=1$ and $(B,\omega_{0})$ is flat. ∎
5.2. The 2-systole on non-rational PSC Kähler surfaces
In this subsection, we study the homological $2$-systole on non-rational PSC Kähler surfaces. Recall that, by the classification of PSC Kähler surfaces, a non-rational PSC Kähler surface is precisely a (possibly blown-up) ruled surface fibred over a curve of genus $g\geq 1$.
By leveraging (5.4), we provide an alternative proof of Theorem 4.5 with an analytic method. It is worth noting that for a Kähler metric the Chern scalar curvature $S(\omega)$ differs from the Riemannian scalar curvature $\text{scal}(g_{\omega})$ by a factor $2$.
Theorem 5.4.
Let $(X,\omega)$ be a non-rational PSC Kähler surface admitting a holomorphic fibration $X\to B$ to a compact Riemann surface $B$ with genus $g(B)\geq 1$. Then
$\min_{X}S_{X}(\omega)\cdot\operatorname{sys}_{2}(\omega)\;\leq\;4\pi.$ (5.6)
Moreover, equality holds if and only if $g(B)=1$, $X$ is covered by $\mathbb{P}^{1}\times\mathbb{C}$ equipped with the product of the Fubini–Study metric on $\mathbb{P}^{1}$ and a flat metric on $\mathbb{C}$, in such a way that $\operatorname{sys}_{2}(\omega)$ is achieved by the $\mathbb{P}^{1}$-fibre.
Proof.
Let $f\colon X\to B$ be the holomorphic fibration, and let $\omega_{0}$ be a constant curvature metric on $B$ of non-positive Gaussian curvature, so that (5.4) holds:
$\int_{B}\left[\int_{D}\left(\frac{\left|\nabla\partial f\right|^{2}}{|\partial f|^{2}}+S_{X}(\omega)\right)\omega\right]\omega_{0}\leq\int_{B}\left(\int_{D}S_{D}(\omega)\,\omega\right)\omega_{0}.$
For every regular value $z$, the fibre $D$ is a smooth rational curve, hence $D\cong\mathbb{P}^{1}$ and $\chi(D)=2$. By Gauss–Bonnet formula,
$\int_{D}S_{D}(\omega)\,\omega=2\pi\,\chi(D)=4\pi.$
Integrating over $B$, we obtain
$4\pi\int_{B}\omega_{0}=2\pi\int_{B}\chi(D)\,\omega_{0}$ $=\int_{B}\left(\int_{D}S_{D}(\omega)\,\omega\right)\omega_{0}$
$\geq\int_{B}\left(\int_{D}S_{X}(\omega)\,\omega\right)\omega_{0}$
$\geq\min_{X}S_{X}(\omega)\cdot\int_{B}\operatorname{Vol}_{\omega}(D)\,\omega_{0}.$
By definition of the homological $2$-systole we have
$\operatorname{Vol}_{\omega}(D)\;\geq\;\operatorname{sys}_{2}(\omega)$
for every regular fibre $D$. Hence
$4\pi\int_{B}\omega_{0}$ $\geq\min_{X}S_{X}(\omega)\cdot\int_{B}\operatorname{Vol}_{\omega}(D)\,\omega_{0}$
$\geq\min_{X}S_{X}(\omega)\cdot\operatorname{sys}_{2}(\omega)\int_{B}\omega_{0},$
and since $\int_{B}\omega_{0}>0$ this yields
$\min_{X}S_{X}(\omega)\cdot\operatorname{sys}_{2}(\omega)\;\leq\;4\pi,$
as claimed. The equality case happens if $B$ is an elliptic curve endowed with a flat metric, $\nabla f$ is parallel, which implies $X$ is isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$, and $(X,\omega)$ is cscK so that $\operatorname{sys}_{2}(\omega)$ is realized by $\mathbb{P}^{1}$-fibre. ∎
Corollary 5.5.
Let $(X,\omega)$ be a non-rational PSC Kähler surface admitting a non-constant holomorphic map $f\colon X\to B$ to a compact Riemann surface $B$ with genus $g(B)\geq 2$. Then
$\min_{X}S_{X}(\omega)\cdot\operatorname{sys}_{2}(\omega)\;<\;4\pi.$
Proof.
If $g(B)\geq 2$, then $B$ is hyperbolic and cannot carry a flat metric. Hence equality in (5.6) cannot occur in Lemma 5.3, nor in Theorem 5.4, and the inequality in Theorem 5.4 is strict. ∎
Example 5.6.
Let $X=\mathbb{P}^{1}\times B$ be a compact complex surface, where $B$ is a compact Riemann surface of genus $g(B)\geq 2$. Equip $X$ with the product Kähler metric
$\omega=\omega_{\mathrm{FS}}\oplus\omega_{B},$
where on $\mathbb{P}^{1}$ we take the Fubini–Study metric normalized by
$\operatorname{Vol}_{\omega_{\mathrm{FS}}}(\mathbb{P}^{1})=\pi,\qquad S_{\mathbb{P}^{1}}(\omega_{\mathrm{FS}})\equiv 4,$
and on $B$ we choose a constant Chern scalar curvature metric with
$S_{B}(\omega_{B})=-4+\varepsilon\quad\text{for some }\varepsilon\in(0,4).$
Then the Chern scalar curvature of the product metric is constant and given by
$S_{X}(\omega)=S_{\mathbb{P}^{1}}(\omega_{\mathrm{FS}})+S_{B}(\omega_{B})=4+(-4+\varepsilon)=\varepsilon,$
so $\min_{X}S_{X}(\omega)=\varepsilon>0$ and $X$ has positive scalar curvature in our convention.
Next we compare the areas of the two basic complex curves:
  • For the $\mathbb{P}^{1}$-fibre $F=\mathbb{P}^{1}\times\{p\}$, calibration by $\omega$ gives
    $\operatorname{Vol}_{\omega}(F)=\operatorname{Vol}_{\omega_{\mathrm{FS}}}(\mathbb{P}^{1})=\pi.$
  • For the $B$-fibre $B_{p}=\{q\}\times B$, Gauss–Bonnet for the Chern scalar curvature gives
    $\int_{B}S_{B}(\omega_{B})\,\omega_{B}=2\pi\chi(B)=2\pi(2-2g(B))=4\pi(1-g(B)).$
    Since $S_{B}(\omega_{B})\equiv-4+\varepsilon<0$, we obtain
    $\operatorname{Vol}_{\omega}(B_{p})=\operatorname{Vol}_{\omega_{B}}(B)=\frac{4\pi(g(B)-1)}{4-\varepsilon}.$
For $g(B)\geq 2$ and $\varepsilon\in(0,4)$ one has $\operatorname{Vol}_{\omega}(B_{p})>\pi$, so the $2$-systole is realized by the $\mathbb{P}^{1}$-fibre:
$\operatorname{sys}_{2}(\omega)=\min\bigl\{\operatorname{Vol}_{\omega}(F),\operatorname{Vol}_{\omega}(B_{p})\bigr\}=\pi.$
Consequently,
$\min_{X}S_{X}(\omega)\cdot\operatorname{sys}_{2}(\omega)=\varepsilon\cdot\pi\;<\;4\pi.$
In particular, this product is independent of the genus $g(B)$, and it can be made arbitrarily close to $4\pi$ by letting $\varepsilon\nearrow 4$.